### v1

# SVD-LLM: Truncation-aware Singular Value Decomposition for Large Language Model Compression

Abstract
The advancements in Large Language Models (LLMs) have been hindered by their substantial sizes, which necessitate LLM compression methods for practical deployment. Singular Value Decomposition (SVD) offers a promising solution for LLM compression. However, state-of-the-art SVD-based LLM compression methods have two key limitations: truncating smaller singular values may lead to higher compression loss, and the lack of update on the remaining model parameters after SVD truncation. In this work, we propose SVD-LLM, a new SVD-based LLM compression method that addresses the limitations of existing methods. SVD-LLM incorporates a truncation-aware data whitening strategy to ensure a direct mapping between singular values and compression loss. Moreover, SVD-LLM adopts a layer-wise closed-form model parameter update strategy to compensate for accuracy degradation caused by SVD truncation. We evaluate SVD-LLM on a total of $11$ datasets and seven models from three different LLM families at four different scales. Our results demonstrate the superiority of SVD-LLM over state-of-the-arts, especially at high model compression ratios. The source code is available at https://github.com/AIoT-MLSys-Lab/SVD-LLM.
1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in a wide range of tasks such as natural language understanding and language generation (Zhao et al., 2023; Gozalo-Brizuela & Garrido-Merchán, 2023). Despite such capabilities, the democratization of LLMs is primarily restricted by their substantial resource demands (Wan et al., 2023; Wang et al., 2024). As an example, GPT-3 (Brown et al., 2020) consumes more than 325 GB of GPU memory to load its weights and requires 5×80GB A100 GPUs to conduct inferences due to its high runtime memory consumption (Xiao et al., 2023).
One of the most effective techniques to reduce the resource demands of LLMs is model compression (Zhu et al., 2023). Compression techniques based on quantization (Frantar et al., 2022; Lin et al., 2023; Xiao et al., 2023), parameter pruning (Ma et al., 2023; Frantar & Alistarh, 2023), and knowledge distillation (Gu et al., 2023; Hsieh et al., 2023) specifically designed for LLMs have been intensively studied. Regardless of their successes, these techniques have their own constraints such as hardware dependency and the need of expensive retraining. Compared to those techniques, compression techniques based on low-rank approximation such as Singular Value Decomposition (SVD) are not limited by those constraints. Moreover, the KV cache of LLMs compressed via SVD at runtime can also be reduced.
Despite these advantages, the potential of SVD for LLM compression has not been thoroughly explored. A few SVD-based LLM compression methods such as ASVD (Yuan et al., 2023) and FWSVD (Hsu et al., 2022) have recently been proposed. However, these methods exhibit severe performance degradation when the model compression ratio is high. Such limitation can be attributed to two fundamental issues invovled in their approaches: ❶ Imprecise Data Preprocessing: although the data preprocessing strategy proposed by ASVD reduces the negative impact of activation outliers, it does not establish a direct relationship between SVD approximation and model compression loss. As a consequence, truncating smaller singular values in SVD could lead to significant compression loss. ❷ Lack of Model Parameter Update after SVD Truncation: as the model compression ratio increases, the number of singular values that need to be truncated in SVD increases as well. To compensate for the accuracy degradation caused by truncating a large number of singular values, updating the remaining model parameters to adapt to the new activation is needed. Unfortunately, existing SVD-based LLM compression methods do not take such update into account, and thus fail to compensate for the accuracy degradation under high model compression ratios.
In this paper, we propose a new SVD-based LLM compression method named SVD-LLM that effectively addresses the two fundamental issues of the existing methods. SVD-LLM differs from existing SVD-based LLM compression methods in two key aspects: ❶ Truncation-Aware Data Whitening: Supported by the theoretical proof, SVD-LLM incorporates a truncation-aware data whitening technique that ensures a direct mapping between singular values and model compression loss. In doing so, the proposed truncation-aware data whitening technique is able to identify which singular values should be truncated to incur minimal model compression loss. ❷ Layer-Wise Closed-Form Model Parameter Update: to compensate for accuracy degradation under high compression ratios, SVD-LLM incorporates a layer-wise closed-form model parameter update strategy to progressively update the remaining weights after SVD truncation layer by layer.
We compare SVD-LLM with three SVD-based methods for LLM compression, including vanilla SVD as well as state-of-the-art methods FWSVD and ASVD. To demonstrate the generability of SVD-LLM, we conduct our evaluation on a total of $11$ datasets and seven models from three different LLM families (LLaMA, Vicuna, and Mistral) at four different scales (7B, 13B, 30B, and 65B). We highlight five of our findings: (1) SVD-LLM consistently outperforms vanilla SVD, FWSVD, and ASVD across all $11$ datasets, three different LLM families, and four different scales. (2) None of the existing SVD-based LLM compression methods report their performance on model compression ratios above 20%. SVD-LLM exhibits significant advantages over vanilla SVD, FWSVD, and ASVD under high compression ratios from 30% to 60%. (3) SVD-LLM also exhibits superiority over state-of-the-art in terms of compression speed. Specifically, when compressing LLaMA-7B under 20% compression ratio on an A100 GPU, ASVD takes about $5.5$ hours whereas SVD-LLM completes the compression process in $15$ minutes. (4) Moreover, SVD-LLM can benefit other LLM compression methods. Our evaluation results show that SVD-LLM is able to further enhance the compression performance of state-of-the-art quantization (GPTQ (Frantar et al., 2022)) and parameter pruning-based (LLM-Pruner (Ma et al., 2023)) LLM compression methods. (5) Lastly, SVD-LLM brings additional benefit beyond compressing the sizes of LLMs, and is also able to reduce the footprint of KV cache during inference at runtime.
2 Related Works
2.1 Large Language Model Compression
LLMs in general contain billion-scale parameters. Applying conventional model compression methods for LLMs is not feasible as they necessitate retraining. To avoid retraining, post-training methods that do not involve retraining LLM in the compression process have been developed. In general, these methods can be grouped into four categories: unstructured pruning, structured pruning, quantization, and low-rank approximation. Specifically, unstructured pruning methods set the individual weights’ elements to zero without changing its shape. A notable contribution is SparseGPT (Frantar & Alistarh, 2023) which prunes the least important weights’ elements with the inversion of the Hessian matrix. However, the irregular sparsification of unstructured pruning is difficult to achieve the desired speedup or memory saving and can only demonstrate its best efficiency on certain hardware architecture such as NVIDIA Ampere GPU. Unlike unstructured pruning, structured pruning methods directly remove the entire channels or other structured components from LLMs, making them easier to implement on hardware. For example, LLM-Pruner (Ma et al., 2023) utilizes a small amount of data to obtain the weight, parameter, and group importance of the coupled structure for pruning with LoRA to recover the precision. However, due to the great modification of the weight matrix in LLM, it suffers from great accuracy degradation, especially under high compression ratios. Quantization methods, on the other hand, achieve model compression by reducing the precision of weight matrices of an LLM. For example, GPTQ (Frantar et al., 2022) employs layer-wise quantization and updates weights with inverse Hessian information. However, quantization has a drawback of only providing a limited range of compression options, typically ranging from 3 to 8 bits. This limited range could prevent the full utilization of the available memory budget.
Figure 1: Overview of SVD-LLM.
2.2 SVD for LLM Compression
Singular Value Decomposition (SVD) is a widely used technique for reducing matrix size by approximating a matrix with two smaller low-ranking matrices (Golub et al., 1987). In the context of LLM compression, only a few SVD-based LLM compression methods have been proposed. Specifically, vanilla SVD only focuses on the compression of the original weight matrix without consider the importance of the parameters, potentially giving a larger compression error. To address this problem, (Hsu et al., 2022) propose FWSVD, which introduces Fisher information to weigh the importance of parameters. However, FWSVD requires complex gradient calculation that demands substantial resources for LLM compression. Another problem of vanilla SVD is the distribution of activation can affect the compression error. To address this issue, (Yuan et al., 2023) propose ASVD, which scales the weight matrix by a diagonal matrix that represents the impact of input channels on weights. However, both FWSVD and ASVD do not establish a direct relationship between singular values and compression errors. As a result, truncating the smallest singular values may not lead to low compression error. Moreover, as the number of truncated singular values increases, the weight matrices in the compressed LLM lose their ability to adjust to new activations. Consequently, these methods incur severe accuracy degradation under high compression ratios.
3 SVD-LLM
Figure 1 provides an overview of the proposed SVD-LLM. SVD-LLM has two main modules: (1) truncation-aware data whitening, and (2) layer-wise closed-form update. As an overview, before compressing LLM, SVD-LLM collects a small random calibration set following the standard procedure (Frantar & Alistarh, 2023; Yuan et al., 2023; Xiao et al., 2023) to provide data for both truncation-aware data whitening and layer-wise closed-form update. For the data whitening process, SVD-LLM obtains the truncation-aware data whitening matrix by calculating the Cholesky decomposition of the multiplication of the input activation and its transpose. After whitening the input activation, SVD-LLM performs SVD and truncates the weight matrices to compress the LLM. Under high model compression ratios, SVD-LLM performs a layer-wise closed-form update to progressively update the remaining weights layer by layer after compression. In the following, we describe the two main modules truncation-aware data whitening, and layer-wise closed-form update in detail.
3.1 Truncation-Aware Data Whitening
Motivation: Due to high variance of the input activation, simply applying vanilla SVD for LLM compression leads to severe accuracy degradation (Yuan et al., 2023). To address this issue, ASVD (Yuan et al., 2023) formulates LLM compression as an optimization problem with the following optimization objective:
$O_{1}$ $=\min(||WX-W^{\prime}X||_{F})$ (1)
where $W$ is the original LLM weight matrix, $X$ is the original input activation of $W$, $W^{\prime}$ is the LLM weight matrix after compression, and $L=||WX-W^{\prime}X||_{F}$ is the compression loss in the form of Frobenious loss.
Specifically, ASVD extracts a diagonal matrix $S_{0}$ from $X$ and inserts $S_{0}S_{0}^{-1}$ into the multiplication of $WX$ to normalize the variance of the input activation, i.e, $WX=(WS_{0})(S_{0}^{-1}X)=\tilde{W_{0}}\tilde{X_{0}}$. Subsequently, SVD is performed on the matrix $\tilde{W_{0}}$ instead of $W$ to obtain the decomposed matrices $\tilde{U_{0}}$, $\tilde{\Sigma_{0}}$, and $\tilde{V_{0}}$. Finally, ASVD applies the truncation of the smallest singular values in $\tilde{\Sigma_{0}}$ to obtain the compressed weight matrix $W^{\prime}_{0}=\tilde{U_{0}}\times\text{Trunc.}(\tilde{\Sigma_{0}})\times\tilde{V_{0}}$.
Although reducing the variance of activation improves the performance, ASVD does not establish a direct relationship between singular values and compression loss (a detailed proof is included in Section A.1). To better illustrate this point, we show two concrete examples in Figure 2(a). In the first example ❶ where one singular value is truncated, truncating the smallest singular value 0.1 results in a higher compression loss (loss $=$ 1.1) than truncating 0.9 (loss $=$ 0.7). In the second example ❷ where multiple singular values are truncated, truncating the smallest two singular values 0.9 and 0.1 also leads to a higher loss (loss $=$ 1.9) than truncating 2.4 and 0.1 (loss $=$ 1.7). Therefore, truncating the smallest singular values does not leads to minimal compression loss.
(a) ASVD.
(b) SVD-LLM.
Figure 2: Comparison of data whitening methods between ASVD and SVD-LLM.
Key Design: The key idea of SVD-LLM is to incorporate a truncation-aware data whitening technique that ensures a direct mapping between singular values and compression loss. To achieve this, SVD-LLM change the whitened activation $\tilde{X}=S^{-1}X$ to be orthogonal so that each channel is independent of each other, i.e., $\tilde{X}\tilde{X}^{T}=S^{-1}XX^{T}(S^{-1})^{T}=I$. Though the matrix $S$ is not unique, we found that the Cholesky decomposition (Meyer, 2023) of $XX^{T}$ is a good solution due to its numerical stability. After that we perform SVD on $\tilde{W}=WS$ instead of the original weight $W$ to obtain three decomposed matrices $\tilde{U},\tilde{\Sigma},\tilde{V}$, where $\tilde{U}=[\tilde{u_{1}},\tilde{u_{2}},\tilde{u_{3}},...,\tilde{u_{r}}]$, $\tilde{\Sigma}=\text{diag}(\tilde{\sigma_{1}},\tilde{\sigma_{2}},\tilde{\sigma_{3}},\cdots,\tilde{\sigma_{r}})$, and $\tilde{V}=[\tilde{v_{1}},\tilde{v_{2}},\tilde{v_{3}},...,\tilde{v_{r}}]$. Lastly, the smallest singular values are truncated to compress $\tilde{\Sigma}$ to $\tilde{\Sigma}_{1}$ and obtain the compressed weight $\tilde{W^{\prime}}=\tilde{U^{\prime}}\tilde{\Sigma}_{1}\tilde{V}^{T}$.
Figure 2(b) illustrates the effect of the proposed truncation-aware data whitening technique. In the first example ❶ where one singular value is truncated, the compression loss is equal to the truncated singular value. In the second example ❷, the compression loss of truncating multiple singular values is equal to the square root of the sum of their squares. As such, under the proposed truncation-aware data whitening technique, truncating the smallest singular values leads to minimal compression loss.
In the following, we provide the theoretical proof on why the proposed truncation-aware data whitening technique ensures a direct mapping between singular values and compression loss under the case of one singular value (Theorem 3.2) and multiple singular values (Corollary 3.3).
Lemma 3.1.
The Frobenius norm of matrix $A$ with dimension $m\times n$ can be deduced into the square root of the trace of its gram matrix, which is:
$\|A\|_{F}$ $\triangleq\left(\sum_{j=1}^{n}\sum_{i=1}^{m}\left|a_{ij}\right|^{2}\right)^{\frac{1}{2}}=\left[\operatorname{trace}\left(A^{T}A\right)\right]^{\frac{1}{2}}$ (2)
Using Lemma 3.1, we obtain the compression loss $L_{i}$ when truncating the $i^{th}$ singular value of $\tilde{W}$ to reduce its rank for compression (the compressed weight is denoted by $\tilde{W}^{\prime}$):
$L_{i}$ $=||(\tilde{W}-\tilde{W}^{\prime})\tilde{X}||_{F}$
$=||\tilde{\sigma}_{i}\tilde{u}_{i}\tilde{v}_{i}^{T}\tilde{X}||_{F}=\tilde{\sigma}_{i}~{}trace(\tilde{u}_{i}\tilde{v}_{i}^{T}\tilde{X}\tilde{X}^{T}\tilde{v}_{i}\tilde{u}_{i}^{T})^{\frac{1}{2}}$ (1)
Since both $\tilde{U}=[\tilde{u}_{1},\tilde{u}_{2},\tilde{u}_{3},...,\tilde{u}_{r}]$ and $\tilde{V}=[\tilde{v}_{1},\tilde{v}_{2},\tilde{v}_{3},...,\tilde{v}_{r}]$ are orthogonal matrices, we have:
[…]
Theorem 3.2.
If the whitening matrix $S$ satisfies $SS^{T}=XX^{T}$, the compression loss $L_{i}$ equals to $\tilde{\sigma_{i}}$.
Proof.
Since $\tilde{X}=S^{-1}X$ and $SS^{T}=XX^{T}$, we can further infer Section 3.1 to get:
$L_{i}$ $=||\tilde{\sigma}_{i}\tilde{u}_{i}\tilde{v}_{i}^{T}\tilde{X}||_{F}=\tilde{\sigma}_{i}~{}trace(\tilde{u}_{i}\tilde{v}_{i}^{T}\tilde{X}\tilde{X}^{T}\tilde{v}_{i}\tilde{u}_{i}^{T})^{\frac{1}{2}}$
$=\tilde{\sigma}_{i}~{}trace(\tilde{u}_{i}\tilde{v}_{i}^{T}S^{-1}XX^{T}(S^{-1})^{T}\tilde{v}_{i}\tilde{u}_{i}^{T})^{\frac{1}{2}}$
$=\tilde{\sigma}_{i}~{}trace(\tilde{u}_{i}\tilde{v}_{i}^{T}S^{-1}SS^{T}(S^{T})^{-1}\tilde{v}_{i}\tilde{u}_{i}^{T})^{\frac{1}{2}}$
$=\tilde{\sigma}_{i}~{}trace(\tilde{u}_{i}\tilde{v}_{i}^{T}\tilde{v}_{i}\tilde{u}_{i}^{T})^{\frac{1}{2}}=\tilde{\sigma}_{i}$
Since $S$ is the Cholesky decomposition of $XX^{T}$ in our algorithm, which satisfies $SS^{T}=XX^{T}$, the loss $L_{i}$ of truncating $\tilde{\sigma_{i}}$ equals to the singular value $\tilde{\sigma_{i}}$ itself. ∎
Corollary 3.3.
If the whitening matrix $S$ satisfies $SS^{T}=XX^{T}$, truncating the smallest singular values leads to the lowest loss $L$ compared to truncating others.
Proof.
Suppose we truncate $\tilde{\sigma}_{m+1},\tilde{\sigma}_{m+2},\tilde{\sigma}_{m+3},...,\tilde{\sigma}_{r}$ in $\tilde{\Sigma}$ for compression, we calculate the square of the loss $L$:
$L^{2}=$ $\left|\left|\sum_{i=m+1}^{r}\tilde{\sigma}_{i}\tilde{u}_{i}\tilde{v}_{i}^{T}\tilde{X}\right|\right|^{2}_{F}$
$=$ $\sum_{j=m+1}^{r}\sum_{i=m+1}^{r}\tilde{\sigma}_{i}\tilde{\sigma}_{j}~{}trace(\tilde{u_{i}}\tilde{v_{i}}^{T}\tilde{X}\tilde{X}^{T}\tilde{v_{j}}\tilde{u_{j}}^{T})$
$=$ $\sum_{i=m+1}^{r}\tilde{\sigma}_{i}^{2}~{}trace(\tilde{u_{i}}\tilde{v_{i}}^{T}\tilde{X}\tilde{X}^{T}\tilde{v_{i}}\tilde{u_{i}}^{T})$
$=$ $\sum_{i=m+1}^{r}(L_{i})^{2}=\sum_{i=1}^{k}(\tilde{\sigma}_{i})^{2}$
The squared loss $L^{2}$ is equal to the sum of the squared singular values. Therefore, truncating the smallest singular values $\tilde{\sigma}_{i}$ achieves the lowest compression loss. ∎
3.2 Layer-Wise Closed-Form Update
Motivation: With the increase of model compression ratio, the weight matrices in the compressed LLM generally lose their ability to adapt to the new activation $X^{\prime}$ derived from the compressed LLM and thus requires the external update for the adaptation during the model compression process. However, existing SVD-based LLM compression techniques do not include this update process, leading to less competitive performance at high compression ratios.
Table 1: Perplexity and Average Adaptation Loss of LLaMA-7B compressed by ASVD on WikiText-2. The ratio of increase in both perplexity and average adaptation loss from the previous compression ratio to the current one is marked in red.
Metric Compression Ratio
20% 30% 40% 50%
Perplexity 11.14 51 (x4.6) 1407 (x28) 15358 (x11)
Average Adaptation Loss 0.83 2.98 (x3.6) 9.63 (x3.2) 44.42 (x4.6)
As listed in Table 1, the perplexity of LLaMA-7B compressed by ASVD increases about 28 times when the compression ratio increases from 30% to 40%. Meanwhile, there is a significant increase in the average of the adaptation loss computed by $L_{a}=||WX^{\prime}-W^{\prime}X^{\prime}||_{F}$, which demonstrates the significant role of minimizing $L_{a}$, especially under high compression ratios.
Key Design: The key idea of SVD-LLM is to incorporate a layer-wise closed-form model parameter update strategy to update the remaining weights after SVD truncation that minimizes the adaptation loss $L_{a}$:
$O_{2}$ $=\min(||WX^{\prime}-W^{\prime}X^{\prime}||_{F})$ (4)
SVD-LLM compresses and updates all the local weights in one decoder layer at a time. In order to perform the update without destroying the low-rank structure of the compressed weights, SVD-LLM only updates the matrix $\tilde{U}$ while keeping $\tilde{\Sigma}_{1}$ (the singular value matrix after truncation) and $\tilde{V}$ fixed:
$\tilde{U^{\prime}}=\arg\min_{\tilde{U^{\prime}}}||WX^{\prime}-\tilde{U^{\prime}}\tilde{\Sigma}_{1}\tilde{V}^{T}X^{\prime}||_{F}$ (5)
To compensate for the error introduced by truncation, we adapt $\tilde{U}$ to its closed-form solution $\tilde{U^{\prime}}=WX^{\prime}X^{\prime T}\tilde{V}\tilde{\Sigma}_{1}(\tilde{\Sigma}_{1}\tilde{V}^{T}X^{\prime}X^{\prime T}\tilde{V}\tilde{\Sigma}_{1})^{-1}$ with the same calibration data used for whitening.
Figure 3: Layer-wise closed-form update in SVD-LLM.
Figure 3 illustrates the overall process. To perform the update in layer $i$, we extract the output $x^{\prime}_{i-1}$ from the previous updated layer $i-1$ as the new input activation, then apply it to every original weight matrix $W$ and its corresponding decomposed matrices $\tilde{\Sigma}_{1}$ and $\tilde{V}$ in layer $i$ to update $\tilde{U}$ based on Equation 5.
After updating $\tilde{U}$ with $\tilde{U^{\prime}}$, we combine and save the truncated matrices $\tilde{U^{\prime}},\tilde{\Sigma}_{1},\tilde{V}$ into two low-rank weight matrices $W_{u}=\tilde{U^{\prime}}(\tilde{\Sigma}_{1})^{1/2}$ and $W_{v}=\tilde{V}(\tilde{\Sigma}_{1})^{1/2}$. Therefore, for the new input activation $X^{\prime}$, its forward pass takes the form $Y=W_{u}W_{v}^{T}X^{\prime}$ in compressed LLM.
4 Experiments
4.1 Experimental Setups
Baselines. We compare SVD-LLM against three baselines including vanilla SVD as well as state-of-the-art SVD-based LLM compression methods FWSVD (Hsu et al., 2022) and ASVD (Yuan et al., 2023).
Models and Datasets. To demonstrate the generability of our method, we evaluate the performance of SVD-LLM and the baselines on seven models from three different LLM families (LLaMA-7B, 13B, 30B, 65B (Touvron et al., 2023a), LLaMA2-7B (Touvron et al., 2023b), Vicuna-7B (Chiang et al., 2023) and Mistral-7B (Jiang et al., 2023) and $11$ datasets including three language modeling datasets (WikiTex-2 (Merity et al., 2016), PTB (Marcus et al., 1993) and C4 (Raffel et al., 2020)) and eight classification datasets (SST-5 (Socher et al., 2013), MNLI (Williams et al., 2018), MRPC (Dolan & Brockett, 2005), COLA (Warstadt et al., 2019), SNLI (Bowman et al., 2015), ARC-e, ARC-c, and ARC-b (Clark et al., 2018)) in zero-shot setting following previous experimental design (Ma et al., 2023).
Implementation Details. To ensure a fair comparison, we followed ASVD (Yuan et al., 2023) to use the calibration set of 256 samples from WikiText-2 for whitening. This calibration set is also used for layer-wise closed-form update. Since layer-wise closed-form update is intended to mitigate accuracy drop under higher compression ratios, we only apply it when compression ratios are at 40% and above. All of our experiments are conducted on Nvidia A100 GPUs.
4.2 Overall Performance
We evaluate the overall performance of SVD-LLM from five aspects: (1) performance on different datasets, (2) performance under high model compression ratios, (3) performance on different LLMs, (4) performance on LLMs with larger scales, and (5) performance with LoRA fine-tuning.
Performance on Different Datasets.
Table 2: Zero-shot performance comparison of LLaMA-7B compressed by SVD-LLM and baselines on three language modeling datasets (measured by perplexity ($\downarrow$)) and eight classification datasets (measured by both individual and average accuracy ($\uparrow$)). The relative performance gain compared to the best-performing baseline is marked in green color inside bracket.
[…]
0% Original 5.68 8.35 7.34 31.54 34.43 66.49 69.13 34.26 26.63 29.49 25.75 39.72
20% SVD 20061 20306 18800 18.91 31.54 33.74 30.87 31.70 20.46 20.00 21.93 26.14
FWSVD 1727 2152 1511 26.38 32.07 42.32 49.09 33.59 23.10 26.44 24.01 32.13
ASVD 11.14 16.55 15.93 26.38 33.50 66.49 68.17 33.09 25.40 23.39 25.64 37.76
SVD-LLM 7.94 ($\downarrow$29%) 16.22 ($\downarrow$2%) 15.84 ($\downarrow$1%) 26.92 35.47 66.55 68.84 34.41 27.16 27.46 26.22 39.13 ($\uparrow$4%)
First, we evaluate the performance of LLaMA-7B compressed by SVD-LLM and the baselines under 20% compression ratio on all $11$ datasets. Table 2 summarizes the results. As shown, SVD-LLM consistently outperforms vanilla SVD, FWSVD and ASVD across all the $11$ datasets. Specifically, on the three language modeling datasets, SVD-LLM outperforms vanilla SVD and FWSVD by a large margin; in comparison with ASVD, SVD-LLM reduces the perplexity by 29%, 2%, and 1% on WikiText-2, PTB, and C4, respectively. On the eight classification datasets, SVD-LLM performs better than the best-performing baseline and achieves 4% higher accuracy on average across the eight datasets.
Performance under High Model Compression Ratios. None of the existing SVD-based LLM compression methods report their performance on model compression ratios above 20%. To bridge this gap, we compare the performance of LLaMA-7B compressed by SVD-LLM and the baselines under compression ratios from 30% to 60%. Table 3 summarizes our results. As shown, SVD-LLM consistently outperforms the three baselines across all the compression ratios. More importantly, compared to low compression ratio scenario in Table 2, SVD-LLM exhibits significant advantages over vanilla SVD, FWSVD, and ASVD under high compression ratios. Specifically, under 30% compression ratio, compared to the best-performing baseline (ASVD), SVD-LLM reduces the perplexity on WikiText-2, PTB, and C4 by 81%, 62%, and 39%, respectively; under compression ratio of 40% and above, SVD-LLM reduces the perplexity by more than 96%. These results indicate that SVD-LLM is more effective in compressing LLMs for more resource-constrained devices such as smartphones and IoT devices.
Table 3: Perplexity ($\downarrow$) of LLaMA-7B compressed by SVD-LLM and baselines under high model compression ratios on WikiText-2, PTB, and C4.
Method 30 % 40% 50% 60%
WikiText-2
SVD 13103 52489 131715 105474
FWSVD 20127 18156 24391 32194
ASVD 51 1407 15358 57057
SVD-LLM 9.56 ($\downarrow$81%) 13.11 ($\downarrow$99%) 23.97 ($\downarrow$99%) 53.74 ($\downarrow$99%)
PTB
SVD 17210 59977 87227 79905
FWSVD 11058 20990 28321 43931
ASVD 70 3292 47690 45218
SVD-LLM 26.39 ($\downarrow$62%) 63.75 ($\downarrow$98%) 150.58 ($\downarrow$99%) 438.58 ($\downarrow$99%)
C4
SVD 20871 47774 79815 106976
FWSVD 7240 12847 23104 29292
ASVD 41 1109 27925 43036
SVD-LLM 25.11 ($\downarrow$39%) 49.83 ($\downarrow$96%) 118.57 ($\downarrow$99%) 345.49 ($\downarrow$99%)
Performance on Different LLMs. To examine the generability of SVD-LLM across different LLMs, we compare the performance between SVD-LLM and the baselines on four models from three different LLM families – LLaMA-7B, LLaMA 2-7B, Mistral-7B, and Vicuna-7B – under 20% compression ratio on WikiText-2. As shown in Table 4, SVD-LLM consistently outperforms vanilla SVD, FWSVD, and ASVD across all four LLMs. In addition, SVD-LLM exhibits more stable performance on different LLM families, especially compared to vanilla SVD and FWSVD.
Table 4: Perplexity ($\downarrow$) of different LLMs under 20% compression ratio on WikiText-2. The relative performance gain compared to the best-performing baseline is marked in green color inside bracket.
Method LLaMA-7B LLaMA 2-7B Mistral-7B Vicuna-7B
SVD 20061 18192 159627 18644
FWSVD 1630 2360 6357 2758
ASVD 11.14 10.10 13.72 16.23
SVD-LLM 7.94 ($\downarrow$29%) 8.50 ($\downarrow$16%) 10.21 ($\downarrow$26%) 6.78 ($\downarrow$58%)
Performance on LLMs with Larger Scales. To examine the generability of SVD-LLM on LLMs across different scales, we compare the performance between SVD-LLM and the baselines on LLaMA series at four different scales – 7B, 13B, 30B, and 65B – under 20% compression ratio on WikiText-2. As shown in Table 5, SVD-LLM consistently outperforms vanilla SVD, FWSVD, and ASVD across all four model sizes. Moreover, both FWSVD and ASVD demand excessive memory resources, causing out of memory (OOM) when compressing LLMs at larger scales even on an A100 GPU due to memory-intensive operations for estimating the importance of weight matrices. In contrast, SVD-LLM does not involve such estimation operations and thus avoids OOM.
Table 5: Perplexity ($\downarrow$) of LLaMA-7B, 13B, 30B, 65B under 20% compression ratio on WikiText-2. Some baselines’ results are not available due to running out of memory (OOM) during model compression. The relative performance gain compared to the best-performing baseline is marked in green color inside bracket.
Method LLaMA-7B LLaMA-13B LLaMA-30B LLaMA-65B
SVD 20061 946.31 54.11 11.27
FWSVD 1630 OOM OOM OOM
ASVD 11.14 6.74 22.71 OOM
SVD-LLM 7.94 ($\downarrow$29%) 6.61 ($\downarrow$2%) 5.63 ($\downarrow$75%) 6.58 ($\downarrow$42%)
Performance with LoRA Fine-Tuning.
Figure 4: Perplexity of LLaMA-7B compressed by SVD-LLM and ASVD (w/ and w/o LoRA) on WikiText-2. The perplexities of ASVD at and beyond $30\%$ compression ratio are higher than $1407$ and thus are not shown in the figure.
LoRA (Hu et al., 2021) is a common fine-tuning technique for LLM. It has been applied with pruning-based LLM compression methods such as LLM-Pruner (Ma et al., 2023) to mitigate accuracy drop after pruning. LoRA can also be combined with SVD-based LLM compression methods by modifying the forward pass of a linear layer as:
$Y$ $=W_{u}^{\prime}W_{v}^{\prime}X$ (6)
where $W_{u}^{\prime}=W_{u}+B_{u}A_{u}$, $W_{v}^{\prime}=W_{v}^{T}+B_{v}A_{v}$, and $A_{u}$, $B_{u}$, $A_{v}$, and $B_{v}$ are low-rank weights fine-tuned using LoRA.
To examine the performance of SVD-LLM in combination with LoRA, we follow the same configuration used in LLM-Pruner (Ma et al., 2023) to fine-tune LLaMA-7B compressed by SVD-LLM and ASVD under compression ratios from 20% to 80% with LoRA. The results are shown in Figure 4. We have three observations. (1) Comparing SVD-LLM with SVD-LLM $+$ LoRA, under the compression ratio between 20% and 50%, the accuracy enhancement brought by LoRA is limited; as the compression ratio increases, LoRA is playing a more significant role in enhancing the performance of SVD-LLM. (2) Compared with ASVD $+$ LoRA, SVD-LLM $+$ LoRA consistently achieves better accuracy across all the compression ratios. In particular, under 70% compression ratio, the perplexity of SVD-LLM $+$ LoRA is 75% lower than ASVD $+$ LoRA. (3) Lastly, even without LoRA, SVD-LLM is able to achieve perplexity comparable to ASVD $+$ LoRA, especially under the compression ratios between 20% and 60%.
4.3 Compression Speed Evaluation
Besides compression performance, we also evaluate the compression speed of SVD-LLM and the baselines. Specifically, we measured the GPU hours used for SVD-LLM and ASVD when compressing LLaMA-7B under 20% compression ratio on an A100 GPU. The results are shown in Table 6. As shown, ASVD takes about $5.5$ hours whereas SVD-LLM completes the compression process in $15$ minutes, which is $32$ times faster. When breaking down the time, most of the time consumed by ASVD is dedicated to calculating the compression ratio of each weight matrix based on its estimated importance through a search process. In contrast, SVD-LLM maintains a consistent compression ratio among all the weight matrices and thus gets rid of the time-consuming search process.
Table 6: Compression speed comparison on LLaMA-7B under 20% compression ratio. The relative speedup is marked in green color inside bracket.
Method SVD-LLM ASVD
Whitening Update Total Preprocessing Search Total
Time 10min 5min 15min ($\downarrow$95%) 5min 5.5h 5.5h
4.4 Ablation Studies
We next conduct ablation studies to evaluate the separate contributions of the two key components (truncation-aware data whitening and layer-wise closed-form update) of SVD-LLM. Let SVD-LLM (W) denote the version of SVD-LLM with truncation-aware data whitening only; SVD-LLM (U) denote the version of SVD-LLM with layer-wise closed-form update only; and SVD-LLM (W+U) denote the version of SVD-LLM with both truncation-aware data whitening and layer-wise closed-form update. The results are shown in Table 7. We have three observations. (1) Both SVD-LLM (W) and SVD-LLM (U) consistently outperform ASVD across all the compression ratios. Notably, when the compression ratio is at and above 40%, both variants reduce the perplexity by more than 99% compared to ASVD. (2) Under 20% and 30% compression ratios, SVD-LLM (W) achieves the lowest perplexity compared to SVD-LLM (U) and SVD-LLM (W+U). (3) Under 40%, 50% and 60% compression ratios, SVD-LLM (W+U) achieves the lowest perplexity compared to SVD-LLM (W) and SVD-LLM (U), highlighting the importance of combining both truncation-aware data whitening and layer-wise closed-form update when compression ratio goes high.
Table 7: Perplexity ($\downarrow$) of LLaMA-7B compressed under different compression ratios on WikiText-2. SVD-LLM (W) denotes the version of SVD-LLM with truncation-aware data whitening only; SVD-LLM (U) denotes the version of SVD-LLM with layer-wise closed-form update only; SVD-LLM (W+U) denotes the version of SVD-LLM with both truncation-aware data whitening and layer-wise closed-form update. The relative performance gain compared to ASVD is marked in green color inside bracket.
Method 20% 30% 40% 50% 60%
ASVD 11.14 51 1407 15358 57057
SVD-LLM (W) 7.94 ($\downarrow$29%) 9.56 ($\downarrow$81%) 13.73 ($\downarrow$99%) 26.11 ($\downarrow$99%) 66.62 ($\downarrow$99%)
SVD-LLM (U) 9.54 ($\downarrow$14%) 12.98 ($\downarrow$75%) 24.16 ($\downarrow$99%) 72.13 ($\downarrow$99%) 204 ($\downarrow$99%)
SVD-LLM (W+U) 8.25 ($\downarrow$26%) 9.95 ($\downarrow$80 %) 13.11 ($\downarrow$99%) 23.97 ($\downarrow$99%) 53.74 ($\downarrow$99%)
SVD-LLM 7.94 ($\downarrow$29%) 9.56 ($\downarrow$81%) 13.11 ($\downarrow$99%) 23.97 ($\downarrow$99%) 53.74 ($\downarrow$99%)
4.5 Benefits to other LLM Compression Methods
SVD-LLM is orthogonal to other LLM compression methods including quantization and parameter pruning. In this experiment, we add the techniques proposed in SVD-LLM into state-of-the-art quantization and parameter pruning-based LLM compression methods to examine how SVD-LLM could further enhance their performance.
Integrate SVD-LLM with Quantization. We select GPTQ (Frantar et al., 2022) as the quantization method. Specifically, we compress LLaMA-7B by GPTQ-4bit combined with SVD-LLM, and compare the compressed model against LLaMA-7B compressed by GPTQ-3bit. As shown in Table 8, combining GPTQ-4bit with SVD-LLM achieves a perplexity that is $18\%$ lower than GPTQ-3bit even with a smaller memory footprint (2.1 GB vs. 2.8 GB). This result demonstrates that compared to directly quantizing using smaller number of bits, GPTQ is able to achieve better compression performance with the help of SVD-LLM.
Table 8: Perplexity ($\downarrow$) of LLaMA-7B compressed by GPTQ w/ and w/o SVD-LLM on WikiText-2. The relative performance gain of combined compression compared to GPTQ-3bit is marked in green color inside bracket.
Method GPTQ-4bit GPTQ-3bit SVD-LLM + GPTQ-4bit
Memory 3.9 GB 2.8 GB 2.1 GB
Perplexity 6.21 16.28 13.29 ($\downarrow$18%)
Integrate SVD-LLM with Parameter Pruning. We select LLM-Pruner (Ma et al., 2023) as the parameter pruning method. Specifically, we compress LLaMA-7B by LLM-Pruner under 30% compression ratio combined with SVD-LLM, and compare the compressed model against LLaMA-7B compressed by LLM-Pruner under 40% compression ratio. As shown in Table 9, LLM-Pruner achieves better compression performance when used in conjunction with SVD-LLM. In particular, with the same memory footprint of 8.8 GB, combining LLM-Pruner under 30% compression ratio with SVD-LLM achieves a perplexity that is $13\%$ lower than LLM-Pruner under 40% compression ratio.
Table 9: Perplexity ($\downarrow$) of LLaMA-7B compressed by LLM-Pruner w/ and w/o SVD-LLM on WikiText-2. The relative performance gain of combined compression compared to LLM-Pruner under 40% compression ratio is marked in green color inside bracket.
Method LLM-Pruner-30% LLM-Pruner-40%
LLM-Pruner-30%
+ SVD-LLM
Memory 9.8 GB 8.8 GB 8.8 GB
Perplexity 9.88 12.21 10.58 ($\downarrow$13%)
4.6 KV Cache Compression with SVD-LLM
SVD-LLM is able to not only compress LLMs but also compress the runtime KV cache at the same time. Specifically, instead of keeping the original intermediate state matrix $m=WX$ with shape $M\times L$ inside the KV cache, after decomposing and compressing $W$ into $W_{u}W_{v}^{T}$, SVD-LLM only needs to store $m^{\prime}=W_{v}^{T}X$ with shape $r\times L$. Therefore, the size of KV cache can be compressed to $\frac{r}{M}$ of the original one. Moreover, since $W_{u}$ is already stored as the weight matrix in the decomposed LLM, the original intermediate state matrix can still be recovered by $m=W_{u}m^{\prime}$ without accuracy drop. Therefore, SVD-LLM provides a unified solution that combines model compression and KV cache compression into a single process. This is different from existing quantization or parameter pruning-based LLM compression methods that need to be combined with other techniques for compressing both weights and KV cache.
In our last experiment, we evaluate this additional benefit on KV cache compression brought by SVD-LLM. This is a new avenue given that KV cache compression has not been evaluated in previous LLM compression studies. Specifically, we measure the peak memory footprint during inference at runtime when generating 128 tokens with batch size of 32 using LLaMA-7B compressed by SVD-LLM under different compression ratios w/ and w/o considering KV cache compression. The results are illustrated in Figure 5 where the difference between blue and yellow bars marked in red represents the reduced footprint of KV cache. As shown, SVD-LLM is able to effectively reduce the footprint of KV cache, and therefore, the peak memory during inference at runtime across all the compression ratios.
Figure 5: Peak memory to generate 128 tokens with batch size of 32 using LLaMA-7B compressed by SVD-LLM under different compression ratios w/ and w/o KV-cache compression. The difference between blue and yellow bars marked in red indicates the reduced footprint of KV cache.
5 Conclusion
In this paper, we presented SVD-LLM, a SVD-based LLM compression method. SVD-LLM proposes a novel truncation-aware data whitening strategy to guide which singular values to be truncated with minimal compression loss. It also introduces a layer-wise closed-form model parameter update scheme to compensate for accuracy degradation under high compression ratios. We have demonstrated the effectiveness of SVD-LLM on 11 datasets and seven models from three LLM families at four scales and have shown its superiority over state-of-the-arts. We also show its effectiveness in further enhancing the performance of other LLM compression methods as well as KV cache reduction.
6 Broader Impacts
We present SVD-LLM, a post-training SVD technique that effectively compresses large language models (LLMs) while maintaining accuracy and robustness. Our method achieves state-of-the-art performance, particularly when dealing with high compression ratios. We demonstrate the value and potential of SVD compression for LLMs by investigating two aspects: (1) the ability of SVD compression to reduce the size of the KV cache, resulting in savings in LLM runtime memory without additionally sacrificing accuracy, and (2) the role of SVD compression in enhancing the compression performance of other LLM compression techniques. Our work contributes to the advancement of compression methods that improve the efficiency and deployability of LLMs for real-world applications.
Appendix A Appendix.
A.1 The compression error of ASVD
The previous state-of-the-art method ASVD introduced a diagonal scaling matrix $S_{0}$ that modifies the weight matrix to reflect the varying significance of different input channels. The linear layer is formulated as $Y=(WS_{0})S_{0}^{-1}X$. The compression is made by keeping the largest $m$ singular value of $WS_{0}$:
$WS_{0}\approx$ $\sum_{i=1}^{m}\sigma^{\prime}_{i}u^{\prime}_{i}{v^{\prime}}_{i}^{T}$
The resulting activation is expressed as:
$Y\approx$ $\sum_{i=1}^{m}\sigma^{\prime}_{i}u^{\prime}_{i}{v^{\prime}}_{i}^{T}S_{0}^{-1}X~{}.$
The compression error $L=||(WS_{0}-\sum_{i=1}^{m}\sigma^{\prime}_{i}u^{\prime}_{i}{v^{\prime}}_{i}^{T})S_{0}^{-1}X||_{F}$ is demonstrated below:
$L^{2}=$ $||(WS_{0}-\sum_{i=1}^{m}\sigma^{\prime}_{i}u^{\prime}_{i}{v^{\prime}}_{i}^{T})S_{0}^{-1}X||_{F}^{2}$
$=$ $\left|\left|\sum_{i=m+1}^{r}\sigma^{\prime}_{i}u^{\prime}_{i}{v^{\prime}}_{i}^{T}S_{0}^{-1}X\right|\right|_{F}^{2}$
$=$ $\sum_{j=m+1}^{r}\sum_{i=m+1}^{r}\sigma^{\prime}_{i}\sigma^{\prime}_{j}~{}trace(u^{\prime}_{i}{v^{\prime}}_{i}^{T}XX^{T}v^{\prime}_{j}{u^{\prime}}_{j}^{T})$
$=$ $\sum_{j=m+1}^{r}\sum_{i=m+1}^{r}\sigma^{\prime}_{i}\sigma^{\prime}_{j}~{}trace({u^{\prime}}_{j}^{T}u^{\prime}_{i}{v^{\prime}}_{i}^{T}S_{0}^{-1}XX^{T}(S_{0}^{-1})^{T}v^{\prime}_{j})$
$=$ $\sum_{i=m+1}^{r}{\sigma^{\prime}}_{i}^{2}~{}trace({v^{\prime}}_{i}^{T}S_{0}^{-1}XX^{T}(S_{0}^{-1})^{T}v^{\prime}_{i})$
$=$ $\sum_{i=m+1}^{r}{\sigma^{\prime}}_{i}^{2}~{}||{v^{\prime}}_{i}^{T}S_{0}^{-1}X||_{F}^{2}~{},$
which is still a complex function that involves the activation $X$, the scaling matrix $S_{0}$, the singular vector $v^{\prime}_{i}$ and the singular value $\sigma^{\prime}_{i}$. As a result compression error is not directly related to the singular value and the conventional SVD compression by truncating the smallest singular values may lead to suboptimal compression error.

### v2

@@
 (4) Moreover, SVD-LLM can benefit other LLM compression methods.
-Our evaluation results show that SVD-LLM is able to further enhance the compression performance of state-of-the-art quantization (GPTQ (Frantar et al., 2022)) and parameter pruning-based (LLM-Pruner (Ma et al., 2023)) LLM compression methods.
+Our evaluation results show that SVD-LLM is able to further enhance the compression performance of well-recognized quantization (GPTQ (Frantar et al., 2022)) and parameter pruning-based (LLM-Pruner (Ma et al., 2023)) LLM compression methods.
 (5) Lastly, SVD-LLM brings additional benefit beyond compressing the sizes of LLMs, and is also able to reduce the footprint of KV cache during inference at runtime.
@@
 To achieve this, SVD-LLM change the whitened activation $\tilde{X}=S^{-1}X$ to be orthogonal so that each channel is independent of each other, i.e., $\tilde{X}\tilde{X}^{T}=S^{-1}XX^{T}(S^{-1})^{T}=I$.
-Though the matrix $S$ is not unique, we found that the Cholesky decomposition (Meyer, 2023) of $XX^{T}$ is a good solution due to its numerical stability.
+Though the matrix $S$ is not unique, we found that the Cholesky decomposition (Meyer, 2000) of $XX^{T}$ is a good solution due to its numerical stability.
 After that we perform SVD on $\tilde{W}=WS$ instead of the original weight $W$ to obtain three decomposed matrices $\tilde{U},\tilde{\Sigma},\tilde{V}$, where $\tilde{U}=[\tilde{u_{1}},\tilde{u_{2}},\tilde{u_{3}},...,\tilde{u_{r}}]$, $\tilde{\Sigma}=\text{diag}(\tilde{\sigma_{1}},\tilde{\sigma_{2}},\tilde{\sigma_{3}},\cdots,\tilde{\sigma_{r}})$, and $\tilde{V}=[\tilde{v_{1}},\tilde{v_{2}},\tilde{v_{3}},...,\tilde{v_{r}}]$.
@@
 Models and Datasets.
-To demonstrate the generability of our method, we evaluate the performance of SVD-LLM and the baselines on seven models from three different LLM families (LLaMA-7B, 13B, 30B, 65B (Touvron et al., 2023a), LLaMA2-7B (Touvron et al., 2023b), Vicuna-7B (Chiang et al., 2023) and Mistral-7B (Jiang et al., 2023) and $11$ datasets including three language modeling datasets (WikiTex-2 (Merity et al., 2016), PTB (Marcus et al., 1993) and C4 (Raffel et al., 2020)) and eight classification datasets (SST-5 (Socher et al., 2013), MNLI (Williams et al., 2018), MRPC (Dolan & Brockett, 2005), COLA (Warstadt et al., 2019), SNLI (Bowman et al., 2015), ARC-e, ARC-c, and ARC-b (Clark et al., 2018)) in zero-shot setting following previous experimental design (Ma et al., 2023).
+To demonstrate the generability of our method, we evaluate the performance of SVD-LLM and the baselines on seven models from three different LLM families (LLaMA-7B, 13B, 30B, 65B (Touvron et al., 2023a), LLaMA2-7B (Touvron et al., 2023b), Vicuna-7B (Chiang et al., 2023) and Mistral-7B (Jiang et al., 2023) and $11$ datasets including three language modeling datasets (WikiTex-2 (Merity et al., 2017), PTB (Marcus et al., 1993) and C4 (Raffel et al., 2020)) and eight classification datasets (SST-5 (Socher et al., 2013), MNLI (Williams et al., 2018), MRPC (Dolan & Brockett, 2005), COLA (Warstadt et al., 2019), SNLI (Bowman et al., 2015), ARC-e, ARC-c, and ARC-b (Clark et al., 2018)) in zero-shot setting following previous experimental design (Ma et al., 2023).
 Implementation Details.
@@
 The perplexities of ASVD at and beyond $30\%$ compression ratio are higher than $1407$ and thus are not shown in the figure.
-LoRA (Hu et al., 2021) is a common fine-tuning technique for LLM.
+LoRA (Hu et al., 2022) is a common fine-tuning technique for LLM.
 It has been applied with pruning-based LLM compression methods such as LLM-Pruner (Ma et al., 2023) to mitigate accuracy drop after pruning.
@@
 Table 6: Compression speed comparison on LLaMA-7B under 20% compression ratio. The relative speedup is marked in green color inside bracket.
-Method SVD-LLM ASVD
+Metric SVD-LLM ASVD
 Whitening Update Total Preprocessing Search Total
@@
 SVD-LLM is orthogonal to other LLM compression methods including quantization and parameter pruning.
-In this experiment, we add the techniques proposed in SVD-LLM into state-of-the-art quantization and parameter pruning-based LLM compression methods to examine how SVD-LLM could further enhance their performance.
+In this experiment, we add the techniques proposed in SVD-LLM into representative quantization and parameter pruning-based LLM compression methods that are widely recognized by the community to examine how SVD-LLM could further enhance their performance.
 Integrate SVD-LLM with Quantization.
@@
 The relative performance gain of combined compression compared to GPTQ-3bit is marked in green color inside bracket.
-Method GPTQ-4bit GPTQ-3bit SVD-LLM + GPTQ-4bit
+Metric GPTQ-4bit GPTQ-3bit SVD-LLM + GPTQ-4bit
 Memory 3.9 GB 2.8 GB 2.1 GB
@@
 The relative performance gain of combined compression compared to LLM-Pruner under 40% compression ratio is marked in green color inside bracket.
-Method LLM-Pruner-30% LLM-Pruner-40%
+Metric LLM-Pruner-30% LLM-Pruner-40%
 LLM-Pruner-30%

### v3

@@
 Singular Value Decomposition (SVD) offers a promising solution for LLM compression.
-However, state-of-the-art SVD-based LLM compression methods have two key limitations: truncating smaller singular values may lead to higher compression loss, and the lack of update on the remaining model parameters after SVD truncation.
+However, state-of-the-art SVD-based LLM compression methods have two key limitations: truncating smaller singular values may lead to higher compression loss, and the lack of update on the compressed weight after SVD truncation.
 In this work, we propose SVD-LLM, a new SVD-based LLM compression method that addresses the limitations of existing methods.
 SVD-LLM incorporates a truncation-aware data whitening strategy to ensure a direct mapping between singular values and compression loss.
-Moreover, SVD-LLM adopts a layer-wise closed-form model parameter update strategy to compensate for accuracy degradation caused by SVD truncation.
-We evaluate SVD-LLM on a total of $11$ datasets and seven models from three different LLM families at four different scales.
+Moreover, SVD-LLM adopts a layer-wise closed-form model parameter update strategy to compensate for accuracy degradation under high compression ratios.
+We evaluate SVD-LLM on a total of $10$ datasets and eight models from three different LLM families at four different scales.
 Our results demonstrate the superiority of SVD-LLM over state-of-the-arts, especially at high model compression ratios.
-The source code is available at https://github.com/AIoT-MLSys-Lab/SVD-LLM.
 1 Introduction
-Large Language Models (LLMs) have demonstrated remarkable capabilities in a wide range of tasks such as natural language understanding and language generation (Zhao et al., 2023; Gozalo-Brizuela & Garrido-Merchán, 2023).
-Despite such capabilities, the democratization of LLMs is primarily restricted by their substantial resource demands (Wan et al., 2023; Wang et al., 2024).
-As an example, GPT-3 (Brown et al., 2020) consumes more than 325 GB of GPU memory to load its weights and requires 5×80GB A100 GPUs to conduct inferences due to its high runtime memory consumption (Xiao et al., 2023).
-One of the most effective techniques to reduce the resource demands of LLMs is model compression (Zhu et al., 2023).
-Compression techniques based on quantization (Frantar et al., 2022; Lin et al., 2023; Xiao et al., 2023), parameter pruning (Ma et al., 2023; Frantar & Alistarh, 2023), and knowledge distillation (Gu et al., 2023; Hsieh et al., 2023) specifically designed for LLMs have been intensively studied.
-Regardless of their successes, these techniques have their own constraints such as hardware dependency and the need of expensive retraining.
-Compared to those techniques, compression techniques based on low-rank approximation such as Singular Value Decomposition (SVD) are not limited by those constraints.
-Moreover, the KV cache of LLMs compressed via SVD at runtime can also be reduced.
+Large Language Models (LLMs) have demonstrated remarkable capabilities in a wide range of tasks such as natural language understanding and language generation [31, 9].
+Despite such capabilities, the democratization of LLMs is primarily restricted by their substantial resource demands [25, 26].
+One of the most effective techniques to reduce the resource demands of LLMs is model compression [32].
+Compression techniques based on quantization [6, 15, 27], parameter pruning [16, 5], and knowledge distillation [10, 11] specifically designed for LLMs have been intensively studied.
+Regardless of their success, these techniques have their own constraints, such as hardware dependency and the need for expensive retraining.
+Compared to those techniques, compression techniques based on low-rank approximation, such as Singular Value Decomposition (SVD) are not limited by those constraints.
 Despite these advantages, the potential of SVD for LLM compression has not been thoroughly explored.
-A few SVD-based LLM compression methods such as ASVD (Yuan et al., 2023) and FWSVD (Hsu et al., 2022) have recently been proposed.
-However, these methods exhibit severe performance degradation when the model compression ratio is high.
-Such limitation can be attributed to two fundamental issues invovled in their approaches: ❶ Imprecise Data Preprocessing: although the data preprocessing strategy proposed by ASVD reduces the negative impact of activation outliers, it does not establish a direct relationship between SVD approximation and model compression loss.
+A few SVD-based LLM compression methods such as ASVD [28] and FWSVD [12] have recently been proposed.
+However, these methods exhibit severe performance degradation when the model compression ratio [^1] The compression ratio refers to the percentage of parameter reduction achieved through compression.
+is high.
+Such limitation can be attributed to two fundamental issues involved in their approaches: ❶ Imprecise Data Preprocessing: although the data preprocessing strategy proposed by ASVD reduces the negative impact of activation outliers, it does not establish a direct relationship between singular values and the model compression loss.
 As a consequence, truncating smaller singular values in SVD could lead to significant compression loss.
 ❷ Lack of Model Parameter Update after SVD Truncation: as the model compression ratio increases, the number of singular values that need to be truncated in SVD increases as well.
-To compensate for the accuracy degradation caused by truncating a large number of singular values, updating the remaining model parameters to adapt to the new activation is needed.
+To compensate for the accuracy degradation caused by truncating a large number of singular values, it is required to update the remaining parameters in the compressed model.
 Unfortunately, existing SVD-based LLM compression methods do not take such update into account, and thus fail to compensate for the accuracy degradation under high model compression ratios.
@@
 In doing so, the proposed truncation-aware data whitening technique is able to identify which singular values should be truncated to incur minimal model compression loss.
-❷ Layer-Wise Closed-Form Model Parameter Update: to compensate for accuracy degradation under high compression ratios, SVD-LLM incorporates a layer-wise closed-form model parameter update strategy to progressively update the remaining weights after SVD truncation layer by layer.
+❷ Layer-Wise Closed-Form Model Parameter Update: to compensate for accuracy degradation under high compression ratios, SVD-LLM incorporates a layer-wise closed-form model parameter update strategy to progressively update the compressed weights layer by layer.
 We compare SVD-LLM with three SVD-based methods for LLM compression, including vanilla SVD as well as state-of-the-art methods FWSVD and ASVD.
-To demonstrate the generability of SVD-LLM, we conduct our evaluation on a total of $11$ datasets and seven models from three different LLM families (LLaMA, Vicuna, and Mistral) at four different scales (7B, 13B, 30B, and 65B).
-We highlight five of our findings: (1) SVD-LLM consistently outperforms vanilla SVD, FWSVD, and ASVD across all $11$ datasets, three different LLM families, and four different scales.
-(2) None of the existing SVD-based LLM compression methods report their performance on model compression ratios above 20%.
-SVD-LLM exhibits significant advantages over vanilla SVD, FWSVD, and ASVD under high compression ratios from 30% to 60%.
-(3) SVD-LLM also exhibits superiority over state-of-the-art in terms of compression speed.
+To demonstrate the generability of SVD-LLM, we conduct our evaluation on a total of $10$ datasets and eight models from three different LLM families (LLaMA, OPT, and Mistral) at four different scales (7B, 13B, 30B, and 65B).
+We highlight six of our findings: (1) SVD-LLM consistently outperforms vanilla SVD, FWSVD, and ASVD across all $10$ datasets, three different LLM families, and four different scales and even exhibits significant advantages under high compression ratios from 30% to 60%.
+(2) SVD-LLM also exhibits superiority over the state-of-the-art in terms of compression speed.
 Specifically, when compressing LLaMA-7B under 20% compression ratio on an A100 GPU, ASVD takes about $5.5$ hours whereas SVD-LLM completes the compression process in $15$ minutes.
-(4) Moreover, SVD-LLM can benefit other LLM compression methods.
-Our evaluation results show that SVD-LLM is able to further enhance the compression performance of well-recognized quantization (GPTQ (Frantar et al., 2022)) and parameter pruning-based (LLM-Pruner (Ma et al., 2023)) LLM compression methods.
+(3) The independent performance of either of the two key components of SVD-LLM still consistently surpasses the performance of the current state-of-the-art SVD compression method under different compression ratios.
+(4) SVD-LLM can benefit other LLM compression methods.
+Our evaluation results show that SVD-LLM is able to further enhance the compression performance of well-recognized quantization (GPTQ [6]) and parameter pruning-based (LLM-Pruner [16]) LLM compression methods.
+(5) SVD-LLM can ensure inference speedup on both GPU and CPU.
+It is able to achieve at most 1.7x speedup on GPU and 1.5x speedup on CPU under the 40% compression ratio.
 (6) Lastly, SVD-LLM brings additional benefit beyond compressing the sizes of LLMs, and is also able to reduce the footprint of KV cache during inference at runtime.
-2 Related Works
-2.1 Large Language Model Compression
-LLMs in general contain billion-scale parameters.
+2 Related Work
+Large Language Model Compression: LLMs in general contain billion-scale parameters.
 Applying conventional model compression methods for LLMs is not feasible as they necessitate retraining.
-To avoid retraining, post-training methods that do not involve retraining LLM in the compression process have been developed.
+To avoid retraining, post-training methods that do not involve retraining LLMs in the compression process have been developed.
 In general, these methods can be grouped into four categories: unstructured pruning, structured pruning, quantization, and low-rank approximation.
 Specifically, unstructured pruning methods set the individual weights’ elements to zero without changing its shape.
-A notable contribution is SparseGPT (Frantar & Alistarh, 2023) which prunes the least important weights’ elements with the inversion of the Hessian matrix.
+A notable contribution is SparseGPT [5] which prunes the least important weight elements with the inversion of the Hessian matrix.
 However, the irregular sparsification of unstructured pruning is difficult to achieve the desired speedup or memory saving and can only demonstrate its best efficiency on certain hardware architecture such as NVIDIA Ampere GPU.
-Unlike unstructured pruning, structured pruning methods directly remove the entire channels or other structured components from LLMs, making them easier to implement on hardware.
-For example, LLM-Pruner (Ma et al., 2023) utilizes a small amount of data to obtain the weight, parameter, and group importance of the coupled structure for pruning with LoRA to recover the precision.
-However, due to the great modification of the weight matrix in LLM, it suffers from great accuracy degradation, especially under high compression ratios.
+Unlike unstructured pruning, structured pruning methods directly remove entire channels or other structured components from LLMs, making them easier to implement on hardware.
+For example, LLM-Pruner [16] utilizes a small amount of data to obtain the weight, parameter, and group importance of the coupled structure for pruning with LoRA to recover precision.
+However, due to the great modification of the weight matrix in LLM, it suffers from a great accuracy degradation, especially under high compression ratios.
 Quantization methods, on the other hand, achieve model compression by reducing the precision of weight matrices of an LLM.
-For example, GPTQ (Frantar et al., 2022) employs layer-wise quantization and updates weights with inverse Hessian information.
-However, quantization has a drawback of only providing a limited range of compression options, typically ranging from 3 to 8 bits.
-This limited range could prevent the full utilization of the available memory budget.
+For example, GPTQ [6] uses layer-wise quantization and updates the weights with inverse Hessian information.
+However, quantization has the drawback of only providing a limited range of compression options, typically ranging from 3 to 8 bits.
+This limited range could prevent full utilization of the available memory budget.
 Figure 1: Overview of SVD-LLM.
-2.2 SVD for LLM Compression
-Singular Value Decomposition (SVD) is a widely used technique for reducing matrix size by approximating a matrix with two smaller low-ranking matrices (Golub et al., 1987).
+SVD for LLM Compression: Singular Value Decomposition (SVD) is a widely used technique to reduce matrix size by approximating a matrix with two smaller low-ranking matrices [8].
 In the context of LLM compression, only a few SVD-based LLM compression methods have been proposed.
-Specifically, vanilla SVD only focuses on the compression of the original weight matrix without consider the importance of the parameters, potentially giving a larger compression error.
-To address this problem, (Hsu et al., 2022) propose FWSVD, which introduces Fisher information to weigh the importance of parameters.
-However, FWSVD requires complex gradient calculation that demands substantial resources for LLM compression.
+Specifically, vanilla SVD only focuses on the compression of the original weight matrix without considering the importance of the parameters, potentially giving a larger compression error.
+To address this problem, [12] propose FWSVD, which introduces Fisher information to weigh the importance of parameters.
+However, FWSVD requires a complex gradient calculation that demands substantial resources for LLM compression.
 Another problem of vanilla SVD is the distribution of activation can affect the compression error.
-To address this issue, (Yuan et al., 2023) propose ASVD, which scales the weight matrix by a diagonal matrix that represents the impact of input channels on weights.
-However, both FWSVD and ASVD do not establish a direct relationship between singular values and compression errors.
-As a result, truncating the smallest singular values may not lead to low compression error.
-Moreover, as the number of truncated singular values increases, the weight matrices in the compressed LLM lose their ability to adjust to new activations.
-Consequently, these methods incur severe accuracy degradation under high compression ratios.
+To address this issue, [28] propose ASVD, which scales the weight matrix by a diagonal matrix that represents the impact of input channels on the weights.
+However, both FWSVD and ASVD do not establish a direct relationship between singular values and compression loss.
+As a result, truncating the smaller singular values may lead to higher compression loss.
+Moreover, as the compression ratio increases, it is necessary to update the compressed weight due to truncating a great number of singular values.
+However, existing methods have no design for this update and thus incur severe accuracy degradation under high compression ratios.
 3 SVD-LLM
-Figure 1 provides an overview of the proposed SVD-LLM.
-SVD-LLM has two main modules: (1) truncation-aware data whitening, and (2) layer-wise closed-form update.
-As an overview, before compressing LLM, SVD-LLM collects a small random calibration set following the standard procedure (Frantar & Alistarh, 2023; Yuan et al., 2023; Xiao et al., 2023) to provide data for both truncation-aware data whitening and layer-wise closed-form update.
-For the data whitening process, SVD-LLM obtains the truncation-aware data whitening matrix by calculating the Cholesky decomposition of the multiplication of the input activation and its transpose.
-After whitening the input activation, SVD-LLM performs SVD and truncates the weight matrices to compress the LLM.
+Figure 1 provides an overview of SVD-LLM.
+At a high level, SVD-LLM is a SVD-based post-training LLM compression method.
+Specifically, following the standard procedure of post-training LLM compression methods [5, 28, 27], SVD-LLM uses a random set of sentences as calibration data to generate activation for truncation-aware data whitening and layer-wise closed-form update for model compression.
+SVD-LLM whitens the activation through Cholesky decomposition, and performs SVD to truncate the weight matrices to compress the LLM.
 Under high model compression ratios, SVD-LLM performs a layer-wise closed-form update to progressively update the remaining weights layer by layer after compression.
-In the following, we describe the two main modules truncation-aware data whitening, and layer-wise closed-form update in detail.
+In the following, we describe both truncation-aware data whitening and layer-wise closed-form update in detail.
+The pseudocode is provided in Section A.2.
 3.1 Truncation-Aware Data Whitening
-Motivation: Due to high variance of the input activation, simply applying vanilla SVD for LLM compression leads to severe accuracy degradation (Yuan et al., 2023).
-To address this issue, ASVD (Yuan et al., 2023) formulates LLM compression as an optimization problem with the following optimization objective:
-$O_{1}$ $=\min(||WX-W^{\prime}X||_{F})$ (1)
-where $W$ is the original LLM weight matrix, $X$ is the original input activation of $W$, $W^{\prime}$ is the LLM weight matrix after compression, and $L=||WX-W^{\prime}X||_{F}$ is the compression loss in the form of Frobenious loss.
-Specifically, ASVD extracts a diagonal matrix $S_{0}$ from $X$ and inserts $S_{0}S_{0}^{-1}$ into the multiplication of $WX$ to normalize the variance of the input activation, i.e, $WX=(WS_{0})(S_{0}^{-1}X)=\tilde{W_{0}}\tilde{X_{0}}$.
-Subsequently, SVD is performed on the matrix $\tilde{W_{0}}$ instead of $W$ to obtain the decomposed matrices $\tilde{U_{0}}$, $\tilde{\Sigma_{0}}$, and $\tilde{V_{0}}$.
-Finally, ASVD applies the truncation of the smallest singular values in $\tilde{\Sigma_{0}}$ to obtain the compressed weight matrix $W^{\prime}_{0}=\tilde{U_{0}}\times\text{Trunc.}(\tilde{\Sigma_{0}})\times\tilde{V_{0}}$.
-Although reducing the variance of activation improves the performance, ASVD does not establish a direct relationship between singular values and compression loss (a detailed proof is included in Section A.1).
-To better illustrate this point, we show two concrete examples in Figure 2(a).
-In the first example ❶ where one singular value is truncated, truncating the smallest singular value 0.1 results in a higher compression loss (loss $=$ 1.1) than truncating 0.9 (loss $=$ 0.7).
+Motivation: Due to high variance of the input activation, simply applying vanilla SVD for LLM compression leads to severe accuracy degradation [28].
+To address this issue, ASVD [28] formulates LLM compression as an optimization problem with the following optimization objective:
+$O=\min(||WX-W^{\prime}X||_{F})$ (1)
+where $W$ is the weight matrix of the original LLM, $X$ is the activation of $W$ given an input, $W^{\prime}$ is the compressed weight matrix, and $L=||WX-W^{\prime}X||_{F}$ is the compression loss in the form of Frobenius loss.
+Specifically, ASVD extracts a diagonal matrix $S_{0}$ from $X$ where each element in the diagonal is the absolute mean value of each channel.
+It then uses $S_{0}$ to normalize $X$ and converts $WX$ into $(WS_{0})(S_{0}^{-1}X)$.
+Subsequently, SVD is performed on $WS_{0}$ to obtain the decomposed matrices $U_{0}$, $\Sigma_{0}$, and $V_{0}$.
+Lastly, ASVD truncates the smallest singular values in $\Sigma_{0}$ to obtain the compressed weight matrix $W^{\prime}_{0}=U_{0}\times\text{Trunc.}(\Sigma_{0})\times V_{0}\times S_{0}^{-1}$.
+Although normalizing the activation improves the performance, ASVD does not establish a direct relationship between singular values and compression loss (a detailed proof is included in Section A.1).
+To better illustrate this point, we show two concrete examples in LABEL:fig:asvd_whitening.
+In the first example ❶ where only one singular value is truncated, truncating the smallest singular value 0.1 results in a higher compression loss (loss $=$ 1.1) than truncating the second smallest singular value 0.9 (loss $=$ 0.7).
 In the second example ❷ where multiple singular values are truncated, truncating the smallest two singular values 0.9 and 0.1 also leads to a higher loss (loss $=$ 1.9) than truncating 2.4 and 0.1 (loss $=$ 1.7).
-Therefore, truncating the smallest singular values does not leads to minimal compression loss.
-(a) ASVD.
-(b) SVD-LLM.
-Figure 2: Comparison of data whitening methods between ASVD and SVD-LLM.
+Hence, truncating the smallest singular values does not lead to minimal loss.
 Key Design: The key idea of SVD-LLM is to incorporate a truncation-aware data whitening technique that ensures a direct mapping between singular values and compression loss.
-To achieve this, SVD-LLM change the whitened activation $\tilde{X}=S^{-1}X$ to be orthogonal so that each channel is independent of each other, i.e., $\tilde{X}\tilde{X}^{T}=S^{-1}XX^{T}(S^{-1})^{T}=I$.
-Though the matrix $S$ is not unique, we found that the Cholesky decomposition (Meyer, 2000) of $XX^{T}$ is a good solution due to its numerical stability.
-After that we perform SVD on $\tilde{W}=WS$ instead of the original weight $W$ to obtain three decomposed matrices $\tilde{U},\tilde{\Sigma},\tilde{V}$, where $\tilde{U}=[\tilde{u_{1}},\tilde{u_{2}},\tilde{u_{3}},...,\tilde{u_{r}}]$, $\tilde{\Sigma}=\text{diag}(\tilde{\sigma_{1}},\tilde{\sigma_{2}},\tilde{\sigma_{3}},\cdots,\tilde{\sigma_{r}})$, and $\tilde{V}=[\tilde{v_{1}},\tilde{v_{2}},\tilde{v_{3}},...,\tilde{v_{r}}]$.
-Lastly, the smallest singular values are truncated to compress $\tilde{\Sigma}$ to $\tilde{\Sigma}_{1}$ and obtain the compressed weight $\tilde{W^{\prime}}=\tilde{U^{\prime}}\tilde{\Sigma}_{1}\tilde{V}^{T}$.
-Figure 2(b) illustrates the effect of the proposed truncation-aware data whitening technique.
-In the first example ❶ where one singular value is truncated, the compression loss is equal to the truncated singular value.
+To achieve this, SVD-LLM enforces the whitened activation $S^{-1}X$ to be orthogonal such that each channel is independent of each other, i.e., $(S^{-1}X)(S^{-1}X)^{T}=S^{-1}XX^{T}(S^{-1})^{T}=I$, where $S$ is derived through Cholesky decomposition [19].
+Then we perform SVD on $WS$ to obtain the decomposed matrices $U,\Sigma,V$, where $U=[u_{1},u_{2},u_{3},...,u_{r}]$, $\Sigma=\text{diag}(\sigma_{1},\sigma_{2},\sigma_{3},\cdots,\sigma_{r})$, and $V=[v_{1},v_{2},v_{3},...,v_{r}]$.
+Lastly, the smallest singular values in $\Sigma$ are truncated to obtain the compressed weight matrix $W^{\prime}=U\times\text{Trunc.}(\Sigma)\times V^{T}\times S^{-1}$.
+LABEL:fig:svdllm_whitening illustrates the effect of the proposed truncation-aware data whitening technique.
+In the first example ❶ where only one singular value is truncated, the compression loss is equal to the truncated singular value.
 In the second example ❷, the compression loss of truncating multiple singular values is equal to the square root of the sum of their squares.
 As such, under the proposed truncation-aware data whitening technique, truncating the smallest singular values leads to minimal compression loss.
-In the following, we provide the theoretical proof on why the proposed truncation-aware data whitening technique ensures a direct mapping between singular values and compression loss under the case of one singular value (Theorem 3.2) and multiple singular values (Corollary 3.3).
+Below, we provide a theoretical proof on why the proposed truncation-aware data whitening technique ensures a direct mapping between singular values and compression loss in the case of one singular value (Theorem 3.2) and multiple singular values (Corollary 3.3).
 Lemma 3.1.
@@
 $\|A\|_{F}\triangleq\left(\sum_{j=1}^{n}\sum_{i=1}^{m}\left|a_{ij}\right|^{2}\right)^{\frac{1}{2}}=\left[\operatorname{trace}\left(A^{T}A\right)\right]^{\frac{1}{2}}$ (2)
-Using Lemma 3.1, we obtain the compression loss $L_{i}$ when truncating the $i^{th}$ singular value of $\tilde{W}$ to reduce its rank for compression (the compressed weight is denoted by $\tilde{W}^{\prime}$):
-$L_{i}$ $=||(\tilde{W}-\tilde{W}^{\prime})\tilde{X}||_{F}$
-$=||\tilde{\sigma}_{i}\tilde{u}_{i}\tilde{v}_{i}^{T}\tilde{X}||_{F}=\tilde{\sigma}_{i}~{}trace(\tilde{u}_{i}\tilde{v}_{i}^{T}\tilde{X}\tilde{X}^{T}\tilde{v}_{i}\tilde{u}_{i}^{T})^{\frac{1}{2}}$ (1)
-Since both $\tilde{U}=[\tilde{u}_{1},\tilde{u}_{2},\tilde{u}_{3},...,\tilde{u}_{r}]$ and $\tilde{V}=[\tilde{v}_{1},\tilde{v}_{2},\tilde{v}_{3},...,\tilde{v}_{r}]$ are orthogonal matrices, we have:
+Using Lemma 3.1, we obtain the compression loss $L_{i}$ when truncating the $i^{th}$ singular value of $S^{-1}X$ to reduce its rank for compression:
+$L_{i}=||(W-W^{\prime})X||_{F}=||\sigma_{i}u_{i}v_{i}^{T}S^{-1}X||_{F}=\sigma_{i}~{}trace(u_{i}v_{i}^{T}S^{-1}XX^{T}(S^{-1})^{T}v_{i}u_{i}^{T})^{\frac{1}{2}}$ (3)
+Since both $U=[u_{1},u_{2},u_{3},...,u_{r}]$ and $V=[v_{1},v_{2},v_{3},...,v_{r}]$ are orthogonal matrices, we have:
+$v_{i}^{T}v_{i}=u_{i}^{T}u_{i}=I;v_{i}^{T}v_{j}=u_{i}^{T}u_{j}=0,\forall i\neq j;trace(v_{i}v_{i}^{T})=trace(u_{i}u_{i}^{T})=1$ (4)
+Theorem 3.2.
+If $S$ is the Cholesky decomposition of $XX^{T}$, the compression loss $L_{i}$ equals to $\sigma_{i}$.
+Proof.
+Since the whitening matrix $S$ is the Cholesky decomposition of $XX^{T}$, we have $SS^{T}=XX^{T}$. We can further infer Equation 3 to obtain:
+$L_{i}=||\sigma_{i}u_{i}v_{i}^{T}S^{-1}X||_{F}=\sigma_{i}~{}trace(u_{i}v_{i}^{T}S^{-1}XX^{T}(S^{-1})^{T}v_{i}u_{i}^{T})^{\frac{1}{2}}=\sigma_{i}~{}trace(u_{i}v_{i}^{T}v_{i}u_{i}^{T})^{\frac{1}{2}}=\sigma_{i}\$ (5)
+Therefore, $L_{i}$ of truncating $\sigma_{i}$ equals to the singular value $\sigma_{i}$ itself. ∎
+Corollary 3.3.
+If $S$ is the Cholesky decomposition of $XX^{T}$, truncating the smallest singular values leads to the lowest loss $L$ compared to truncating others.
+Proof.
+If we truncate $\sigma_{m+1},\sigma_{m+2},\sigma_{m+3},...,\sigma_{r}$ in $\Sigma$ for compression, the square of the loss $L$ is:
+$\begin{split}L^{2}=&\left|\left|\sum_{i=m+1}^{r}\sigma_{i}u_{i}v_{i}^{T}S^{-1}X\right|\right|^{2}_{F}=\sum_{j=m+1}^{r}\sum_{i=m+1}^{r}\sigma_{i}\sigma_{j}~{}trace(u_{i}v_{i}^{T}S^{-1}XX^{T}(S^{-1})^{T}v_{j}u_{j}^{T})\\ =&\sum_{i=m+1}^{r}\sigma_{i}^{2}~{}trace(u_{i}v_{i}^{T}S^{-1}XX^{T}(S^{-1})^{T}v_{i}u_{i}^{T})=\sum_{i=m+1}^{r}(L_{i})^{2}=\sum_{i=1}^{k}(\sigma_{i})^{2}\end{split}$ (6)
+The squared loss $L^{2}$ is equal to the sum of the squared singular values. Therefore, truncating the smallest singular values $\sigma_{i}$ achieves the lowest compression loss. ∎
+3.2 Layer-Wise Closed-Form Update
+Motivation: Given the same calibration data as input, the compressed weight matrix $W^{\prime}$ generates a new activation $X^{\prime}$ that is different from $X$ generated by the original weight matrix $W$.
+As the compression ratio increases, $W$ needs to truncate a larger number of singular values to obtain $W^{\prime}$, thus $X^{\prime}$ deviates further from $X$.
+Therefore, it becomes necessary to design a strategy for updating $W^{\prime}$ to minimize $||WX^{\prime}-W^{\prime}X^{\prime}||_{F}$.
+However, existing SVD-based LLM compression methods have no design of parameter update after compression, leading to less competitive performance at high compression ratios.
+Figure 3: Layer-Wise Closed-Form Update.
 […]
-Theorem 3.2.
-If the whitening matrix $S$ satisfies $SS^{T}=XX^{T}$, the compression loss $L_{i}$ equals to $\tilde{\sigma_{i}}$.
-Proof.
-Since $\tilde{X}=S^{-1}X$ and $SS^{T}=XX^{T}$, we can further infer Section 3.1 to get:
-$L_{i}$ $=||\tilde{\sigma}_{i}\tilde{u}_{i}\tilde{v}_{i}^{T}\tilde{X}||_{F}=\tilde{\sigma}_{i}~{}trace(\tilde{u}_{i}\tilde{v}_{i}^{T}\tilde{X}\tilde{X}^{T}\tilde{v}_{i}\tilde{u}_{i}^{T})^{\frac{1}{2}}$
-$=\tilde{\sigma}_{i}~{}trace(\tilde{u}_{i}\tilde{v}_{i}^{T}S^{-1}XX^{T}(S^{-1})^{T}\tilde{v}_{i}\tilde{u}_{i}^{T})^{\frac{1}{2}}$
-$=\tilde{\sigma}_{i}~{}trace(\tilde{u}_{i}\tilde{v}_{i}^{T}S^{-1}SS^{T}(S^{T})^{-1}\tilde{v}_{i}\tilde{u}_{i}^{T})^{\frac{1}{2}}$
-$=\tilde{\sigma}_{i}~{}trace(\tilde{u}_{i}\tilde{v}_{i}^{T}\tilde{v}_{i}\tilde{u}_{i}^{T})^{\frac{1}{2}}=\tilde{\sigma}_{i}$
-Since $S$ is the Cholesky decomposition of $XX^{T}$ in our algorithm, which satisfies $SS^{T}=XX^{T}$, the loss $L_{i}$ of truncating $\tilde{\sigma_{i}}$ equals to the singular value $\tilde{\sigma_{i}}$ itself.
-∎
-Corollary 3.3.
-If the whitening matrix $S$ satisfies $SS^{T}=XX^{T}$, truncating the smallest singular values leads to the lowest loss $L$ compared to truncating others.
-Proof.
-Suppose we truncate $\tilde{\sigma}_{m+1},\tilde{\sigma}_{m+2},\tilde{\sigma}_{m+3},...,\tilde{\sigma}_{r}$ in $\tilde{\Sigma}$ for compression, we calculate the square of the loss $L$:
-$L^{2}=$ $\left|\left|\sum_{i=m+1}^{r}\tilde{\sigma}_{i}\tilde{u}_{i}\tilde{v}_{i}^{T}\tilde{X}\right|\right|^{2}_{F}$
-$=$ $\sum_{j=m+1}^{r}\sum_{i=m+1}^{r}\tilde{\sigma}_{i}\tilde{\sigma}_{j}~{}trace(\tilde{u_{i}}\tilde{v_{i}}^{T}\tilde{X}\tilde{X}^{T}\tilde{v_{j}}\tilde{u_{j}}^{T})$
-$=$ $\sum_{i=m+1}^{r}\tilde{\sigma}_{i}^{2}~{}trace(\tilde{u_{i}}\tilde{v_{i}}^{T}\tilde{X}\tilde{X}^{T}\tilde{v_{i}}\tilde{u_{i}}^{T})$
-$=$ $\sum_{i=m+1}^{r}(L_{i})^{2}=\sum_{i=1}^{k}(\tilde{\sigma}_{i})^{2}$
-The squared loss $L^{2}$ is equal to the sum of the squared singular values. Therefore, truncating the smallest singular values $\tilde{\sigma}_{i}$ achieves the lowest compression loss. ∎
-3.2 Layer-Wise Closed-Form Update
-Motivation: With the increase of model compression ratio, the weight matrices in the compressed LLM generally lose their ability to adapt to the new activation $X^{\prime}$ derived from the compressed LLM and thus requires the external update for the adaptation during the model compression process.
-However, existing SVD-based LLM compression techniques do not include this update process, leading to less competitive performance at high compression ratios.
-Table 1: Perplexity and Average Adaptation Loss of LLaMA-7B compressed by ASVD on WikiText-2.
-The ratio of increase in both perplexity and average adaptation loss from the previous compression ratio to the current one is marked in red.
-Metric Compression Ratio
-20% 30% 40% 50%
-Perplexity 11.14 51 (x4.6) 1407 (x28) 15358 (x11)
-Average Adaptation Loss 0.83 2.98 (x3.6) 9.63 (x3.2) 44.42 (x4.6)
-As listed in Table 1, the perplexity of LLaMA-7B compressed by ASVD increases about 28 times when the compression ratio increases from 30% to 40%.
-Meanwhile, there is a significant increase in the average of the adaptation loss computed by $L_{a}=||WX^{\prime}-W^{\prime}X^{\prime}||_{F}$, which demonstrates the significant role of minimizing $L_{a}$, especially under high compression ratios.
-Key Design: The key idea of SVD-LLM is to incorporate a layer-wise closed-form model parameter update strategy to update the remaining weights after SVD truncation that minimizes the adaptation loss $L_{a}$:
-$O_{2}$ $=\min(||WX^{\prime}-W^{\prime}X^{\prime}||_{F})$ (4)
-SVD-LLM compresses and updates all the local weights in one decoder layer at a time.
-In order to perform the update without destroying the low-rank structure of the compressed weights, SVD-LLM only updates the matrix $\tilde{U}$ while keeping $\tilde{\Sigma}_{1}$ (the singular value matrix after truncation) and $\tilde{V}$ fixed:
-$\tilde{U^{\prime}}=\arg\min_{\tilde{U^{\prime}}}||WX^{\prime}-\tilde{U^{\prime}}\tilde{\Sigma}_{1}\tilde{V}^{T}X^{\prime}||_{F}$ (5)
-To compensate for the error introduced by truncation, we adapt $\tilde{U}$ to its closed-form solution $\tilde{U^{\prime}}=WX^{\prime}X^{\prime T}\tilde{V}\tilde{\Sigma}_{1}(\tilde{\Sigma}_{1}\tilde{V}^{T}X^{\prime}X^{\prime T}\tilde{V}\tilde{\Sigma}_{1})^{-1}$ with the same calibration data used for whitening.
-Figure 3: Layer-wise closed-form update in SVD-LLM.
-Figure 3 illustrates the overall process.
-To perform the update in layer $i$, we extract the output $x^{\prime}_{i-1}$ from the previous updated layer $i-1$ as the new input activation, then apply it to every original weight matrix $W$ and its corresponding decomposed matrices $\tilde{\Sigma}_{1}$ and $\tilde{V}$ in layer $i$ to update $\tilde{U}$ based on Equation 5.
-After updating $\tilde{U}$ with $\tilde{U^{\prime}}$, we combine and save the truncated matrices $\tilde{U^{\prime}},\tilde{\Sigma}_{1},\tilde{V}$ into two low-rank weight matrices $W_{u}=\tilde{U^{\prime}}(\tilde{\Sigma}_{1})^{1/2}$ and $W_{v}=\tilde{V}(\tilde{\Sigma}_{1})^{1/2}$.
-Therefore, for the new input activation $X^{\prime}$, its forward pass takes the form $Y=W_{u}W_{v}^{T}X^{\prime}$ in compressed LLM.
+$\begin{split}U_{i}^{\prime}&=\arg\min_{U_{i}^{\prime}}||W_{i}X^{\prime}_{i-1}-W^{\prime}_{i}X^{\prime}_{i-1}||_{F}\\ &=W_{i}X^{\prime}_{i-1}D^{T}(DD^{T})^{-1},D=\text{Trunc.}(\Sigma)_{i}V_{i}^{T}S_{i}^{-1}X^{\prime}_{i-1}\end{split}$ (7)
 4 Experiments
-4.1 Experimental Setups
-Baselines. We compare SVD-LLM against three baselines including vanilla SVD as well as state-of-the-art SVD-based LLM compression methods FWSVD (Hsu et al., 2022) and ASVD (Yuan et al., 2023).
+Baselines. We compare SVD-LLM against three baselines including vanilla SVD as well as state-of-the-art SVD-based LLM compression methods FWSVD [12] and ASVD [28].
 Models and Datasets.
-To demonstrate the generability of our method, we evaluate the performance of SVD-LLM and the baselines on seven models from three different LLM families (LLaMA-7B, 13B, 30B, 65B (Touvron et al., 2023a), LLaMA2-7B (Touvron et al., 2023b), Vicuna-7B (Chiang et al., 2023) and Mistral-7B (Jiang et al., 2023) and $11$ datasets including three language modeling datasets (WikiTex-2 (Merity et al., 2017), PTB (Marcus et al., 1993) and C4 (Raffel et al., 2020)) and eight classification datasets (SST-5 (Socher et al., 2013), MNLI (Williams et al., 2018), MRPC (Dolan & Brockett, 2005), COLA (Warstadt et al., 2019), SNLI (Bowman et al., 2015), ARC-e, ARC-c, and ARC-b (Clark et al., 2018)) in zero-shot setting following previous experimental design (Ma et al., 2023).
+To demonstrate the generability of our method, we evaluate the performance of SVD-LLM and the baselines on eight models from three different LLM families (LLaMA-7B, 13B, 30B, 65B [23], LLaMA2-7B [24], OPT-6.7B [30], Vicuna-7B [3] and Mistral-7B [14]) and $10$ datasets including three language modeling datasets (WikiText-2 [18], PTB [17] and C4 [21]) and seven classification datasets (OpenbookQA [20], WinoGrande [22], HellaSwag [29], PIQA [2], MathQA [1], ARC-e, and ARC-c [4]) in zero-shot setting with LM-Evaluation-Harness framework [7].
 Implementation Details.
-To ensure a fair comparison, we followed ASVD (Yuan et al., 2023) to use the calibration set of 256 samples from WikiText-2 for whitening.
-This calibration set is also used for layer-wise closed-form update.
-Since layer-wise closed-form update is intended to mitigate accuracy drop under higher compression ratios, we only apply it when compression ratios are at 40% and above.
+To ensure a fair comparison, we followed ASVD [28] to randomly select 256 samples from WikiText-2 as the calibration data.
+Since layer-wise closed-form update is intended to mitigate the accuracy drop under higher compression ratios, we only apply it when the compression ratios are at 40% and above.
 All of our experiments are conducted on Nvidia A100 GPUs.
 4.1 Overall Performance
-We evaluate the overall performance of SVD-LLM from five aspects: (1) performance on different datasets, (2) performance under high model compression ratios, (3) performance on different LLMs, (4) performance on LLMs with larger scales, and (5) performance with LoRA fine-tuning.
-Performance on Different Datasets.
-Table 2: Zero-shot performance comparison of LLaMA-7B compressed by SVD-LLM and baselines on three language modeling datasets (measured by perplexity ($\downarrow$)) and eight classification datasets (measured by both individual and average accuracy ($\uparrow$)).
+We evaluate the overall performance of SVD-LLM from four aspects: (1) performance under different compression ratios, (2) performance on different LLMs, (3) performance on LLMs with larger scales, and (4) performance with LoRA fine-tuning (See Section A.3).
+Some generated contents by the compressed LLM are listed in Section A.4 to provide a more straightforward comparison.
+Performance under Different Compression Ratios.
+Table 1: Zero-shot performance of LLaMA-7B compressed by SVD-LLM and baselines under 20% to 60% compression ratio on three language modeling datasets (measured by perplexity ($\downarrow$)) and seven common sense reasoning datasets (measured by both individual and average accuracy ($\uparrow$)).
+The best performance is marked in bold.
 The relative performance gain compared to the best-performing baseline is marked in green color inside bracket.
 […]
-0% Original 5.68 8.35 7.34 31.54 34.43 66.49 69.13 34.26 26.63 29.49 25.75 39.72
-20% SVD 20061 20306 18800 18.91 31.54 33.74 30.87 31.70 20.46 20.00 21.93 26.14
-FWSVD 1727 2152 1511 26.38 32.07 42.32 49.09 33.59 23.10 26.44 24.01 32.13
-ASVD 11.14 16.55 15.93 26.38 33.50 66.49 68.17 33.09 25.40 23.39 25.64 37.76
-SVD-LLM 7.94 ($\downarrow$29%) 16.22 ($\downarrow$2%) 15.84 ($\downarrow$1%) 26.92 35.47 66.55 68.84 34.41 27.16 27.46 26.22 39.13 ($\uparrow$4%)
-First, we evaluate the performance of LLaMA-7B compressed by SVD-LLM and the baselines under 20% compression ratio on all $11$ datasets.
-Table 2 summarizes the results.
-As shown, SVD-LLM consistently outperforms vanilla SVD, FWSVD and ASVD across all the $11$ datasets.
-Specifically, on the three language modeling datasets, SVD-LLM outperforms vanilla SVD and FWSVD by a large margin; in comparison with ASVD, SVD-LLM reduces the perplexity by 29%, 2%, and 1% on WikiText-2, PTB, and C4, respectively.
-On the eight classification datasets, SVD-LLM performs better than the best-performing baseline and achieves 4% higher accuracy on average across the eight datasets.
-Performance under High Model Compression Ratios.
-None of the existing SVD-based LLM compression methods report their performance on model compression ratios above 20%.
-To bridge this gap, we compare the performance of LLaMA-7B compressed by SVD-LLM and the baselines under compression ratios from 30% to 60%.
-Table 3 summarizes our results.
-As shown, SVD-LLM consistently outperforms the three baselines across all the compression ratios.
-More importantly, compared to low compression ratio scenario in Table 2, SVD-LLM exhibits significant advantages over vanilla SVD, FWSVD, and ASVD under high compression ratios.
-Specifically, under 30% compression ratio, compared to the best-performing baseline (ASVD), SVD-LLM reduces the perplexity on WikiText-2, PTB, and C4 by 81%, 62%, and 39%, respectively; under compression ratio of 40% and above, SVD-LLM reduces the perplexity by more than 96%.
+0% Original 5.68 8.35 7.34 0.28 0.67 0.67 0.56 0.38 0.78 0.27 0.52
+20% SVD 20061 20306 18800 0.14 0.27 0.51 0.26 0.21 0.53 0.21 0.31
+FWSVD 1727 2152 1511 0.15 0.31 0.50 0.26 0.23 0.56 0.21 0.32
+ASVD 11.14 16.55 15.93 0.25 0.53 0.64 0.41 0.27 0.68 0.24 0.43
+SVD-LLM 7.94 ($\downarrow$29%) 16.22 ($\downarrow$2%) 15.84 ($\downarrow$1%) 0.22 0.58 0.63 0.43 0.29 0.69 0.24 0.44 ($\uparrow$2%)
+30% SVD 13103 17210 20871 0.13 0.26 0.51 0.26 0.21 0.54 0.22 0.30
+FWSVD 20127 11058 7240 0.17 0.26 0.49 0.26 0.22 0.51 0.19 0.30
+ASVD 51 70 41 0.18 0.43 0.53 0.37 0.25 0.65 0.21 0.38
+SVD-LLM 9.56 ($\downarrow$81%) 26.39 ($\downarrow$62%) 25.11 ($\downarrow$39%) 0.20 0.48 0.59 0.37 0.26 0.65 0.22 0.40 ($\uparrow$5%)
+40% SVD 52489 59977 47774 0.15 0.26 0.52 0.26 0.22 0.53 0.20 0.30
+FWSVD 18156 20990 12847 0.16 0.26 0.51 0.26 0.22 0.53 0.21 0.30
+ASVD 1407 3292 1109 0.13 0.28 0.48 0.26 0.22 0.55 0.19 0.30
+SVD-LLM 13.11 ($\downarrow$99%) 63.75 ($\downarrow$98%) 49.83 ($\downarrow$96%) 0.19 0.42 0.58 0.33 0.25 0.60 0.21 0.37 ($\uparrow$23%)
+50% SVD 131715 87227 79815 0.16 0.26 0.50 0.26 0.23 0.52 0.19 0.30
+FWSVD 24391 28321 23104 0.12 0.26 0.50 0.26 0.23 0.53 0.20 0.30
+ASVD 15358 47690 27925 0.12 0.26 0.51 0.26 0.22 0.52 0.19 0.30
+SVD-LLM 23.97 ($\downarrow$99%) 150.58 ($\downarrow$99%) 118.57 ($\downarrow$99%) 0.16 0.33 0.54 0.29 0.23 0.56 0.21 0.33 ($\uparrow$10%)
+60% SVD 105474 79905 106976 0.16 0.26 0.50 0.26 0.22 0.52 0.21 0.30
+FWSVD 32194 43931 29292 0.15 0.26 0.49 0.26 0.22 0.53 0.18 0.30
+ASVD 57057 45218 43036 0.12 0.26 0.49 0.26 0.21 0.51 0.18 0.29
+SVD-LLM 53.74 ($\downarrow$99%) 438.58 ($\downarrow$99%) 345.49 ($\downarrow$99%) 0.14 0.28 0.50 0.27 0.22 0.55 0.21 0.31 ($\uparrow$7%)
+First, we evaluate the performance of LLaMA-7B compressed by SVD-LLM and the baselines under 20% to 60% compression ratios.
+Section 4.1 summarizes the results on all $10$ datasets.
+As shown, SVD-LLM consistently outperforms vanilla SVD, FWSVD and ASVD across all of the compression ratios.
+More importantly, compared to the low compression ratio scenario in Section 4.1, SVD-LLM exhibits significant advantages over vanilla SVD, FWSVD, and ASVD under high compression ratios.
+Specifically, under 30% compression ratio, compared to the best-performing baseline (ASVD), SVD-LLM reduces the perplexity on WikiText-2, PTB, and C4 by 81%, 62%, and 39%, respectively; When the compression ratio reaches 40% and above, SVD-LLM reduces the perplexity by more than 96%.
 These results indicate that SVD-LLM is more effective in compressing LLMs for more resource-constrained devices such as smartphones and IoT devices.
-Table 3: Perplexity ($\downarrow$) of LLaMA-7B compressed by SVD-LLM and baselines under high model compression ratios on WikiText-2, PTB, and C4.
-Method 30 % 40% 50% 60%
-WikiText-2
-SVD 13103 52489 131715 105474
-FWSVD 20127 18156 24391 32194
-ASVD 51 1407 15358 57057
-SVD-LLM 9.56 ($\downarrow$81%) 13.11 ($\downarrow$99%) 23.97 ($\downarrow$99%) 53.74 ($\downarrow$99%)
-PTB
-SVD 17210 59977 87227 79905
-FWSVD 11058 20990 28321 43931
-ASVD 70 3292 47690 45218
-SVD-LLM 26.39 ($\downarrow$62%) 63.75 ($\downarrow$98%) 150.58 ($\downarrow$99%) 438.58 ($\downarrow$99%)
-C4
-SVD 20871 47774 79815 106976
-FWSVD 7240 12847 23104 29292
-ASVD 41 1109 27925 43036
-SVD-LLM 25.11 ($\downarrow$39%) 49.83 ($\downarrow$96%) 118.57 ($\downarrow$99%) 345.49 ($\downarrow$99%)
+On the seven classification datasets, SVD-LLM performs better than the best-performing baseline on most of the datasets and consistently achieves at least 2% higher average accuracy across all the compression ratios.
 Performance on Different LLMs.
-To examine the generability of SVD-LLM across different LLMs, we compare the performance between SVD-LLM and the baselines on four models from three different LLM families – LLaMA-7B, LLaMA 2-7B, Mistral-7B, and Vicuna-7B – under 20% compression ratio on WikiText-2.
-As shown in Table 4, SVD-LLM consistently outperforms vanilla SVD, FWSVD, and ASVD across all four LLMs.
+To examine the generability of SVD-LLM across different LLMs, we compare the performance between SVD-LLM and the baselines on four different models, including OPT-6.7B, LLaMA 2-7B, Mistral-7B, and Vicuna-7B under 20% compression ratio on WikiText-2 and the common sense reasoning datasets (See Section A.5).
+The result on WikiText-2 is shown in Section 4.1, SVD-LLM consistently outperforms vanilla SVD, FWSVD, and ASVD across all four LLMs.
 In addition, SVD-LLM exhibits more stable performance on different LLM families, especially compared to vanilla SVD and FWSVD.
-Table 4: Perplexity ($\downarrow$) of different LLMs under 20% compression ratio on WikiText-2.
+Table 2: Perplexity ($\downarrow$) of four different LLMs including OPT-6.7B, LLaMA 2-7B, Mistral-7B, and Vicuna-7B under 20% compression ratio on WikiText-2.
 The relative performance gain compared to the best-performing baseline is marked in green color inside bracket.
-Method LLaMA-7B LLaMA 2-7B Mistral-7B Vicuna-7B
-SVD 20061 18192 159627 18644
-FWSVD 1630 2360 6357 2758
-ASVD 11.14 10.10 13.72 16.23
-SVD-LLM 7.94 ($\downarrow$29%) 8.50 ($\downarrow$16%) 10.21 ($\downarrow$26%) 6.78 ($\downarrow$58%)
-Performance on LLMs with Larger Scales.
-To examine the generability of SVD-LLM on LLMs across different scales, we compare the performance between SVD-LLM and the baselines on LLaMA series at four different scales – 7B, 13B, 30B, and 65B – under 20% compression ratio on WikiText-2.
-As shown in Table 5, SVD-LLM consistently outperforms vanilla SVD, FWSVD, and ASVD across all four model sizes.
-Moreover, both FWSVD and ASVD demand excessive memory resources, causing out of memory (OOM) when compressing LLMs at larger scales even on an A100 GPU due to memory-intensive operations for estimating the importance of weight matrices.
-In contrast, SVD-LLM does not involve such estimation operations and thus avoids OOM.
+Method OPT-6.7B LLaMA 2-7B Mistral-7B Vicuna-7B
+SVD 66275 18192 159627 18644
+FWSVD 14559 2360 6357 2758
+ASVD 82 10.10 13.72 16.23
+SVD-LLM 16.04 ($\downarrow$80%) 8.50 ($\downarrow$16%) 10.21 ($\downarrow$26%) 6.78 ($\downarrow$58%)
 Table 3: Perplexity ($\downarrow$) of LLaMA-7B, 13B, 30B, 65B under 20% compression ratio on WikiText-2.
@@
 SVD-LLM 7.94 ($\downarrow$29%) 6.61 ($\downarrow$2%) 5.63 ($\downarrow$75%) 6.58 ($\downarrow$42%)
-Performance with LoRA Fine-Tuning.
-Figure 4: Perplexity of LLaMA-7B compressed by SVD-LLM and ASVD (w/ and w/o LoRA) on WikiText-2.
-The perplexities of ASVD at and beyond $30\%$ compression ratio are higher than $1407$ and thus are not shown in the figure.
-LoRA (Hu et al., 2022) is a common fine-tuning technique for LLM.
-It has been applied with pruning-based LLM compression methods such as LLM-Pruner (Ma et al., 2023) to mitigate accuracy drop after pruning.
-LoRA can also be combined with SVD-based LLM compression methods by modifying the forward pass of a linear layer as:
-$Y$ $=W_{u}^{\prime}W_{v}^{\prime}X$ (6)
-where $W_{u}^{\prime}=W_{u}+B_{u}A_{u}$, $W_{v}^{\prime}=W_{v}^{T}+B_{v}A_{v}$, and $A_{u}$, $B_{u}$, $A_{v}$, and $B_{v}$ are low-rank weights fine-tuned using LoRA.
-To examine the performance of SVD-LLM in combination with LoRA, we follow the same configuration used in LLM-Pruner (Ma et al., 2023) to fine-tune LLaMA-7B compressed by SVD-LLM and ASVD under compression ratios from 20% to 80% with LoRA.
-The results are shown in Figure 4.
-We have three observations.
-(1) Comparing SVD-LLM with SVD-LLM $+$ LoRA, under the compression ratio between 20% and 50%, the accuracy enhancement brought by LoRA is limited; as the compression ratio increases, LoRA is playing a more significant role in enhancing the performance of SVD-LLM.
-(2) Compared with ASVD $+$ LoRA, SVD-LLM $+$ LoRA consistently achieves better accuracy across all the compression ratios.
-In particular, under 70% compression ratio, the perplexity of SVD-LLM $+$ LoRA is 75% lower than ASVD $+$ LoRA.
-(3) Lastly, even without LoRA, SVD-LLM is able to achieve perplexity comparable to ASVD $+$ LoRA, especially under the compression ratios between 20% and 60%.
+Performance on LLMs with Larger Scales.
+To examine the generability of SVD-LLM on LLMs across different scales, we compare the performance between SVD-LLM and the baselines on LLaMA series at four different scales – 7B, 13B, 30B, and 65B – under 20% compression ratio on WikiText-2.
+As shown in Section 4.1, SVD-LLM consistently outperforms vanilla SVD, FWSVD, and ASVD across all four model sizes.
+Moreover, both FWSVD and ASVD demand excessive memory resources, causing out of memory (OOM) when compressing LLMs at larger scales even on an A100 GPU due to memory-intensive operations for estimating the importance of weight matrices.
+In contrast, SVD-LLM does not involve such estimation operations and thus avoids OOM.
 4.2 Compression Speed Evaluation
 Besides compression performance, we also evaluate the compression speed of SVD-LLM and the baselines.
-Specifically, we measured the GPU hours used for SVD-LLM and ASVD when compressing LLaMA-7B under 20% compression ratio on an A100 GPU.
-The results are shown in Table 6.
+Specifically, we measured the GPU hours used for SVD-LLM and ASVD when compressing LLaMA-7B under the 20% compression ratio on an A100 GPU.
+The results are shown in Section 4.2.
 As shown, ASVD takes about $5.5$ hours whereas SVD-LLM completes the compression process in $15$ minutes, which is $32$ times faster.
 When breaking down the time, most of the time consumed by ASVD is dedicated to calculating the compression ratio of each weight matrix based on its estimated importance through a search process.
-In contrast, SVD-LLM maintains a consistent compression ratio among all the weight matrices and thus gets rid of the time-consuming search process.
-Table 6: Compression speed comparison on LLaMA-7B under 20% compression ratio. The relative speedup is marked in green color inside bracket.
+In contrast, SVD-LLM maintains a consistent compression ratio among all weight matrices and thus gets rid of the time-consuming search process.
+Table 4: Compression time of SVD-LLM and ASVD on LLaMA-7B under 20% compression ratio. The relative speedup is marked in green color inside bracket.
 Metric SVD-LLM ASVD
-Whitening Update Total Preprocessing Search Total
+White Update Total Normalize Search Total
 Time 10min 5min 15min ($\downarrow$95%) 5min 5.5h 5.5h
-4.4 Ablation Studies
-We next conduct ablation studies to evaluate the separate contributions of the two key components (truncation-aware data whitening and layer-wise closed-form update) of SVD-LLM.
-Let SVD-LLM (W) denote the version of SVD-LLM with truncation-aware data whitening only; SVD-LLM (U) denote the version of SVD-LLM with layer-wise closed-form update only; and SVD-LLM (W+U) denote the version of SVD-LLM with both truncation-aware data whitening and layer-wise closed-form update.
-The results are shown in Table 7.
-We have three observations.
-(1) Both SVD-LLM (W) and SVD-LLM (U) consistently outperform ASVD across all the compression ratios.
-Notably, when the compression ratio is at and above 40%, both variants reduce the perplexity by more than 99% compared to ASVD.
-(2) Under 20% and 30% compression ratios, SVD-LLM (W) achieves the lowest perplexity compared to SVD-LLM (U) and SVD-LLM (W+U).
-(3) Under 40%, 50% and 60% compression ratios, SVD-LLM (W+U) achieves the lowest perplexity compared to SVD-LLM (W) and SVD-LLM (U), highlighting the importance of combining both truncation-aware data whitening and layer-wise closed-form update when compression ratio goes high.
-Table 7: Perplexity ($\downarrow$) of LLaMA-7B compressed under different compression ratios on WikiText-2.
+Table 5: Performance of LLaMA-7B compressed by SVD-LLM under 20% and 30% compression ratios using calibration data randomly sampled from WikiText-2 (by default in our paper) and C4.
+The performance on WikiText-2, PTB, and C4 is reported by perplexity ($\downarrow$), while the performance on OpenbookQA and HellaSwag are reported by accuracy ($\uparrow$).
+The relative performance drop (gain) for data sampled from C4 compared to that sampled from WikiText-2 is marked in red (green) color inside bracket.
+Ratio WikiText-2 PTB C4 Openb. HellaS.
+Calibration data sampled from WikiText-2 (seed=3)
+20% 7.94 16.22 15.84 0.22 0.43
+30% 9.56 26.39 25.11 0.20 0.37
+Calibration data sampled from C4 (seed=3)
+20% 8.62($\uparrow$9%) 14.63($\downarrow$9%) 13.77($\downarrow$13%) 0.23($\uparrow$5%) 0.43
+30% 10.67($\uparrow$12%) 25.07($\downarrow$5%) 22.37($\downarrow$10%) 0.22($\uparrow$9%) 0.42($\uparrow$14%)
+4.3 Ablation Study
+Table 6: Perplexity ($\downarrow$) of compressed LLaMA-7B on WikiText-2.
 SVD-LLM (W) denotes the version of SVD-LLM with truncation-aware data whitening only; SVD-LLM (U) denotes the version of SVD-LLM with layer-wise closed-form update only; SVD-LLM (W+U) denotes the version of SVD-LLM with both truncation-aware data whitening and layer-wise closed-form update.
-The relative performance gain compared to ASVD is marked in green color inside bracket.
+The relative performance gain compared to ASVD is marked in green color.
 Method 20% 30% 40% 50% 60%
@@
 SVD-LLM 7.94 ($\downarrow$29%) 9.56 ($\downarrow$81%) 13.11 ($\downarrow$99%) 23.97 ($\downarrow$99%) 53.74 ($\downarrow$99%)
+Modular Sensitivity Study: We conduct ablation studies to evaluate the separate contributions of the two key components (truncation-aware data whitening and layer-wise closed-form update) of SVD-LLM.
+Let SVD-LLM (W) denote the version of SVD-LLM with truncation-aware data whitening only; SVD-LLM (U) denote the version of SVD-LLM with layer-wise closed-form update only; and SVD-LLM (W+U) denote the version of SVD-LLM with both truncation-aware data whitening and layer-wise closed-form update.
+The results are shown in Section 4.3.
+We have three observations.
+(1) Both SVD-LLM (W) and SVD-LLM (U) consistently outperform ASVD across all the compression ratios.
+Notably, when the compression ratio is at and above 40%, both variants reduce the perplexity by more than 99% compared to ASVD.
+(2) Under 20% and 30% compression ratios, SVD-LLM (W) achieves the lowest perplexity compared to SVD-LLM (U) and SVD-LLM (W+U).
+(3) Under 40%, 50% and 60% compression ratios, SVD-LLM (W+U) achieves the lowest perplexity compared to SVD-LLM (W) and SVD-LLM (U), highlighting the importance of combining both truncation-aware data whitening and layer-wise closed-form update when compression ratio goes high.
+Calibration Data Analysis: We next analyze the impact of calibration data used for both truncation-aware data whitening and layer-wise closed-form update on the compression performance.
+LABEL:fig:cali and Section 4.2 summarize the performance of compressed LLaMA-7B when changing three key characteristics of the calibration data, the number of the calibration data, the seed used to randomly sample the calibration data, and the data set from which the calibration data is sampled.
+As shown, changing any of the three characteristics only causes a tiny disturbance of less than 15% to the final performance, demonstrating that SVD-LLM is less sensitive to the design of the calibration data to compress LLM.
 4.4 Benefits to other LLM Compression Methods
 SVD-LLM is orthogonal to other LLM compression methods including quantization and parameter pruning.
-In this experiment, we add the techniques proposed in SVD-LLM into representative quantization and parameter pruning-based LLM compression methods that are widely recognized by the community to examine how SVD-LLM could further enhance their performance.
+In this experiment, we combine SVD-LLM with quantization and parameter pruning-based LLM compression methods that are widely recognized by the community to examine how SVD-LLM could further enhance their performance.
 Integrate SVD-LLM with Quantization.
-We select GPTQ (Frantar et al., 2022) as the quantization method.
+We select GPTQ [6] as the quantization method.
 Specifically, we compress LLaMA-7B by GPTQ-4bit combined with SVD-LLM, and compare the compressed model against LLaMA-7B compressed by GPTQ-3bit.
-As shown in Table 8, combining GPTQ-4bit with SVD-LLM achieves a perplexity that is $18\%$ lower than GPTQ-3bit even with a smaller memory footprint (2.1 GB vs.
+As shown in Section 4.4, combining GPTQ-4bit with SVD-LLM achieves a perplexity that is $18\%$ lower than GPTQ-3bit even with a smaller memory footprint (2.1 GB vs.
 2.8 GB).
-This result demonstrates that compared to directly quantizing using smaller number of bits, GPTQ is able to achieve better compression performance with the help of SVD-LLM.
+This result demonstrates that compared to directly quantizing using smaller number of bits, GPTQ achieves better compression performance with the help of SVD-LLM.
+Integrate SVD-LLM with Parameter Pruning.
+We select LLM-Pruner [16] as the parameter pruning method.
+Specifically, we compress LLaMA-7B by LLM-Pruner under 30% compression ratio combined with SVD-LLM, and compare the compressed model against LLaMA-7B compressed by LLM-Pruner under 40% compression ratio.
+As shown in Table 4.4, LLM-Pruner achieves better compression performance when used in conjunction with SVD-LLM.
+In particular, with the same memory footprint of 8.8 GB, combining LLM-Pruner under 30% compression ratio with SVD-LLM achieves a perplexity that is $13\%$ lower than LLM-Pruner under 40% compression ratio.
 Table 7: Perplexity ($\downarrow$) of LLaMA-7B compressed by GPTQ w/ and w/o SVD-LLM on WikiText-2.
@@
 Perplexity 6.21 16.28 13.29 ($\downarrow$18%)
-Integrate SVD-LLM with Parameter Pruning.
-We select LLM-Pruner (Ma et al., 2023) as the parameter pruning method.
-Specifically, we compress LLaMA-7B by LLM-Pruner under 30% compression ratio combined with SVD-LLM, and compare the compressed model against LLaMA-7B compressed by LLM-Pruner under 40% compression ratio.
-As shown in Table 9, LLM-Pruner achieves better compression performance when used in conjunction with SVD-LLM.
-In particular, with the same memory footprint of 8.8 GB, combining LLM-Pruner under 30% compression ratio with SVD-LLM achieves a perplexity that is $13\%$ lower than LLM-Pruner under 40% compression ratio.
 Table 8: Perplexity ($\downarrow$) of LLaMA-7B compressed by LLM-Pruner w/ and w/o SVD-LLM on WikiText-2.
-The relative performance gain of combined compression compared to LLM-Pruner under 40% compression ratio is marked in green color inside bracket.
+The relative performance gain of combined compression compared to LLM-Pruner under 40% compression ratio is marked in green color.
 Metric LLM-Pruner-30% LLM-Pruner-40%
@@
 Perplexity 9.88 12.21 10.58 ($\downarrow$13%)
-4.6 KV Cache Compression with SVD-LLM
+4.5 Benefits of Inference Speedup
+SVD-LLM is capable to achieve inference speedup.
+To demonstrate this advantage, we measure the number of tokens that the original LLaMA-7B and its compressed version by SVD-LLM can generate on average per second with different batch size and sequence length.
+LABEL:fig:gpu_throughput show the results on the GPU and CPU.
+As shown, SVD-LLM consistently ensures an acceleration in the generation speed across all the compression ratios illustrated in the figure.
+More importantly, this enhancement becomes more significant as the batch size increases and the sequence length decreases, resulting in a maximum speedup of 1.7x on GPU and 1.5x on CPU under the 40% compression ratio, where the model performance remains acceptable according to Section 4.1.
+These results highlight the effectiveness of SVD-LLM in improving the efficiency of LLM for real-world usage.
+Figure 10: Peak memory to generate 128 tokens with batch size of 32 using LLaMA-7B compressed by SVD-LLM under different compression ratios w/ and w/o KV-cache compression.
+The difference between the blue and yellow bars marked in red indicates the reduced footprint of the KV cache.
+4.6 Benefits of KV Cache Compression
 SVD-LLM is able to not only compress LLMs but also compress the runtime KV cache at the same time.
 Specifically, instead of keeping the original intermediate state matrix $m=WX$ with shape $M\times L$ inside the KV cache, after decomposing and compressing $W$ into $W_{u}W_{v}^{T}$, SVD-LLM only needs to store $m^{\prime}=W_{v}^{T}X$ with shape $r\times L$.
-Therefore, the size of KV cache can be compressed to $\frac{r}{M}$ of the original one.
+Therefore, the size of the KV cache can be compressed to $\frac{r}{M}$ of the original.
 Moreover, since $W_{u}$ is already stored as the weight matrix in the decomposed LLM, the original intermediate state matrix can still be recovered by $m=W_{u}m^{\prime}$ without accuracy drop.
 Therefore, SVD-LLM provides a unified solution that combines model compression and KV cache compression into a single process.
-This is different from existing quantization or parameter pruning-based LLM compression methods that need to be combined with other techniques for compressing both weights and KV cache.
-In our last experiment, we evaluate this additional benefit on KV cache compression brought by SVD-LLM.
-This is a new avenue given that KV cache compression has not been evaluated in previous LLM compression studies.
-Specifically, we measure the peak memory footprint during inference at runtime when generating 128 tokens with batch size of 32 using LLaMA-7B compressed by SVD-LLM under different compression ratios w/ and w/o considering KV cache compression.
-The results are illustrated in Figure 5 where the difference between blue and yellow bars marked in red represents the reduced footprint of KV cache.
-As shown, SVD-LLM is able to effectively reduce the footprint of KV cache, and therefore, the peak memory during inference at runtime across all the compression ratios.
-Figure 5: Peak memory to generate 128 tokens with batch size of 32 using LLaMA-7B compressed by SVD-LLM under different compression ratios w/ and w/o KV-cache compression.
-The difference between blue and yellow bars marked in red indicates the reduced footprint of KV cache.
+This is different from existing quantization or parameter pruning-based LLM compression methods that need to be combined with other techniques for compressing both weights and KV cache..
+In our last experiment, we evaluate this benefit on KV cache compression brought by SVD-LLM.
+This is a new avenue since KV cache compression has not been evaluated in previous LLM compression studies.
+Specifically, we measure the peak memory footprint during inference when generating 128 tokens with batch size of 32 using LLaMA-7B compressed by SVD-LLM under different compression ratios w/ and w/o considering KV cache compression.
+The results are illustrated in Figure 10 where the difference between the blue and yellow bars marked in red represents the reduced footprint of the KV cache.
+As shown, SVD-LLM is able to effectively reduce the footprint of KV cache.
+Therefore, the peak memory during inference at runtime across all the compression ratios.
 5 Conclusion
@@
 It also introduces a layer-wise closed-form model parameter update scheme to compensate for accuracy degradation under high compression ratios.
-We have demonstrated the effectiveness of SVD-LLM on 11 datasets and seven models from three LLM families at four scales and have shown its superiority over state-of-the-arts.
-We also show its effectiveness in further enhancing the performance of other LLM compression methods as well as KV cache reduction.
-6 Broader Impacts
-We present SVD-LLM, a post-training SVD technique that effectively compresses large language models (LLMs) while maintaining accuracy and robustness.
-Our method achieves state-of-the-art performance, particularly when dealing with high compression ratios.
-We demonstrate the value and potential of SVD compression for LLMs by investigating two aspects: (1) the ability of SVD compression to reduce the size of the KV cache, resulting in savings in LLM runtime memory without additionally sacrificing accuracy, and (2) the role of SVD compression in enhancing the compression performance of other LLM compression techniques.
-Our work contributes to the advancement of compression methods that improve the efficiency and deployability of LLMs for real-world applications.
+We have demonstrated the effectiveness of SVD-LLM on $10$ datasets and seven models from three LLM families at four scales and have shown its superiority over state-of-the-arts.
+We also show its effectiveness in further enhancing the performance of other LLM compression methods.
 Appendix A Appendix.
@@
 $=$ $\sum_{i=m+1}^{r}{\sigma^{\prime}}_{i}^{2}~{}||{v^{\prime}}_{i}^{T}S_{0}^{-1}X||_{F}^{2}~{},$
-which is still a complex function that involves the activation $X$, the scaling matrix $S_{0}$, the singular vector $v^{\prime}_{i}$ and the singular value $\sigma^{\prime}_{i}$.
-As a result compression error is not directly related to the singular value and the conventional SVD compression by truncating the smallest singular values may lead to suboptimal compression error.
+which is still a complex function that involves the activation $X$, the diagonal matrix $S_{0}$, the singular vector $v^{\prime}_{i}$ and the singular value $\sigma^{\prime}_{i}$.
+As a result, compression error is not directly related to the singular value, and the conventional SVD compression by truncating the smallest singular values may lead to suboptimal compression error.
+A.2 Pseudocode for SVD-LLM
+Algorithm 1 shows the pseudocode of SVD-LLM.
+Before compression, SVD-LLM randomly collects a small amount of sentences as the calibration data $C$, it then runs the truncation-aware data whitening process as shown in Algorithm 2 to obtain the set of whitening matrix $\text{Set}_{S}$ for the weight to compress.
+After that, it runs the SVD and truncation with $\text{Set}_{S}$ on each weight matrix in the LLM.
+Instead of directly finishing the whole compression, it stores the decomposed matrices and further utilizes these matrices to run the layer-wise closed-form update as shown in Algorithm 3.
+Algorithm 1 Pseudocode for SVD-LLM
+1:Input: $M$: Original LLM
+2:Output: $M^{\prime}$: Compressed LLM by SVD-LLM
+3:procedure SVD-LLM($M$)
+4:     Randomly collect several sentences as the calibration data $C$
+5:     $\text{Set}_{S}\leftarrow\textsc{Truncation-Aware Data Whitening}(M,C)$
+6:     $\text{Set}_{SVD}\leftarrow\emptyset$triangleright Initialize the set of decomposed matrices for the weight to compress
+7:     $\text{Set}_{W}\leftarrow M$triangleright Obtain the set of weights in $M$ to compress
+8:     for $W$ in $\text{Set}_{W}$ do
+9:         $S\leftarrow\text{Set}_{S}(W)$triangleright Extract the whitening matrix of current weight $W$
+10:         $U,\Sigma,V\leftarrow\text{SVD}(WS)$triangleright Apply singular value decomposition on $W$
+11:         $\Sigma_{1}\leftarrow\text{Trunc.}(\Sigma)$triangleright Truncate the smallest singular values in $\Sigma$
+12:         $\text{Set}_{SVD}\leftarrow(U,\Sigma_{1},V)\cap\text{Set}_{SVD}$triangleright Store the decomposed matrices of the weight in the set
+13:     end for
+14:     $M^{\prime}\leftarrow\textsc{Layer-Wise Closed-Form Update}(M,C,\text{Set}_{S},\text{Set}_{SVD})$
+15:     return $M^{\prime}$
+16:end procedure
+Algorithm 2 Pseudocode for Truncation-Aware Data Whitening
+1:Input: $M$: Original LLM
+2:Input: $C$: Calibration Data
+3:Output: $\text{Set}_{S}$: Set of whitening matrices for the weight to compress in $M$
+4:procedure Truncation-Aware Data Whitening($M,C$)
+5:     $\text{Set}_{S}\leftarrow\emptyset$triangleright Initialize the set of whitening matrices
+6:     $\text{Set}_{W}\leftarrow M$triangleright Obtain the set of weights in $M$ to compress
+7:     for $W$ in $\text{Set}_{W}$ do
+8:         $X\leftarrow M(W,D)$triangleright Obtain the input activation of the weight matrix $W$
+9:         $S\leftarrow\text{Cholesky\_Decomposition}(XX^{T})$triangleright Apply cholesky decomposition on $XX^{T}$
+10:         $\text{Set}_{S}\leftarrow S\cup\text{Set}_{S}$triangleright Store the whitening weight matrix in the set
+11:     end for
+12:     return $\text{Set}_{S}$
+13:end procedure
+Algorithm 3 Pseudocode for Layer-Wise Closed-Form Update
+1:Input: $M$: Original LLM
+2:Input: $C$: Calibration Data
+3:Input: $\text{Set}_{S}$: Set of whitening matrices for the weight to compress in $M$
+4:Input: $\text{Set}_{SVD}$: Set of decomposed matrices for the weight to compress in $M$
+5:Output: $M^{\prime}$: Compressed LLM by SVD-LLM
+6:procedure Layer-Wise Closed-Form Update($M,C,\text{Set}_{S},\text{Set}_{SVD}$)
+7:     $M^{\prime}\leftarrow M$triangleright Initialize $M^{\prime}$ with $M$
+8:     $\text{Set}_{L}\leftarrow M^{\prime}$triangleright Obtain the set of encoder and decoder layers in $M^{\prime}$
+9:     $X^{\prime}\leftarrow M^{\prime}(C)$triangleright Obtain the input activation of the first layer in $M^{\prime}$
+10:     for $L$ in $\text{Set}_{L}$ do
+11:         $\text{Set}_{W}\leftarrow L$triangleright Obtain the set of weights in $L$ to compress
+12:         for $W$ in $\text{Set}_{W}$ do
+13:              $S\leftarrow\text{Set}_{S}(W)$triangleright Extract the whitening matrix of current weight $W$
+14:              $U,\Sigma_{1},V\leftarrow\text{Set}_{SVD}(W)$triangleright Obtain the decomposed matrices of $W$ from $\text{Set}_{SVD}$
+15:              $U^{\prime}=WX^{\prime}X^{\prime T}(S^{-1})^{T}V\Sigma_{1}(\Sigma_{1}V^{T}S^{-1}X^{\prime}X^{\prime T}(S^{-1})^{T}V\Sigma_{1})^{-1}$triangleright Closed-form
+16:              $W_{u}\leftarrow U^{\prime}(\Sigma_{1})^{1/2},W_{v}\leftarrow S^{-1}V(\Sigma_{1})^{1/2}$triangleright Obtain two low-rank matrices
+17:              $L(W)\leftarrow L(W_{u},W_{v})$triangleright Replace $W$ with $W_{u}$ and $W_{v}$ in $L$
+18:         end for
+19:         $X^{\prime}\leftarrow L(X^{\prime})$triangleright Use the compressed layer to calculate the new input activation $X^{\prime}$ for the next layer
+20:     end for
+21:     return $M^{\prime}$
+22:end procedure
+Table 9: Perplexity of LLaMA-7B compressed by SVD-LLM and ASVD (w/ and w/o LoRA) on WikiText-2.
+Method 20% 30% 40% 50% 60%
+ASVD 11.14 51 1407 15358 57057
+ASVD + LoRA 7.37 10.16 14.86 21.83 44.81
+SVD-LLM 7.94 9.56 13.11 23.97 53.74
+SVD-LLM + LoRA 8.28 9.14 10.65 13.26 17.93
+A.3 Performance with LoRA Fine-Tuning.
+LoRA [13] is a common fine-tuning technique for LLM.
+It has been applied with pruning-based LLM compression methods such as LLM-Pruner [16] to mitigate accuracy drop after pruning.
+LoRA can also be combined with SVD-based LLM compression methods by modifying the forward pass of a linear layer as:
+$Y$ $=W_{u}^{\prime}W_{v}^{\prime}X$ (8)
+where $W_{u}^{\prime}=W_{u}+B_{u}A_{u}$, $W_{v}^{\prime}=W_{v}^{T}+B_{v}A_{v}$, and $A_{u}$, $B_{u}$, $A_{v}$, and $B_{v}$ are low-rank weights fine-tuned using LoRA.
+To examine the performance of SVD-LLM in combination with LoRA, we follow the same configuration used in LLM-Pruner [16] to fine-tune LLaMA-7B compressed by SVD-LLM and ASVD under the compression ratios from 20% to 80% with LoRA.
+The results are shown in Section A.2.
+We have three observations.
+(1) Comparing SVD-LLM with SVD-LLM $+$ LoRA, under the compression ratio between 20% and 50%, the accuracy enhancement brought by LoRA is limited; as the compression ratio increases, LoRA plays a more significant role in improving the performance of SVD-LLM.
+(2) Compared to ASVD $+$ LoRA, SVD-LLM $+$ LoRA consistently achieves better accuracy across all the compression ratios.
+In particular, under 70% compression ratio, the perplexity of SVD-LLM $+$ LoRA is 75% lower than that of ASVD $+$ LoRA.
+(3) Finally, even without LoRA, SVD-LLM is able to achieve perplexity comparable to ASVD $+$ LoRA, especially under the compression ratios between 20% and 60%.
+A.4 Generated Content from the Compressed Model
+We also compare some examples of sentences generated by LLaMA-7B compressed with SVD-LLM and ASVD in Table 10.
+As shown, the sentences generated by the model compressed by SVD-LLM exhibit better fluency, relevance, and informativeness compared to that compressed by ASVD.
+More importantly, when the compression ratio is increased to 30%, the previous state-of-the-art method ASVD completely loses its generation ability.
+In contrast, even when the compression ratio is up to 40%, SVD-LLM is still capable of generating complete sentences.
+Table 10: An example of contents generated by the compressed LLaMA-7B by SVD-LLM and and under different compression ratios.
+The input is marked in bold and the normal texts are the generated sentences.
+Ratio SVD-LLM ASVD
+Original Building a website can be done in 10 simple steps: Step 1: Register an domain name; Step 2: Get a hosting package; Step 3: Develop a website; Step 4: Design the website; Step 5: Make the website live; Step 6: Perform on-page SEO; Step 7: Promote the website; Step 8: Evaluate the results; Step 9: Improve the website; Step 10: Build a website repeat …
+20% Building a website can be done in 10 simple steps: 1) Identifying an appropriate theme.
+2 ) Determining which types of Web site would be of interest and which could be implemented with minimal overhead and resources.
+3 ) Identify the audience 4 ) Choose Web site software.
+5 ) Select Web site developers.
+6 ) Design Web site 7 ) Identify Web site technology 8 ) Design Web site layout 9 ) Define Web site aesthetics and 10 ) Define Web site access controls … Building a website can be done in 10 simple steps: 1, Build up your website.
+You will need to select a name for your website.
+As a beginner, it is a good idea to select your company name or product name as your website name.
+You can use your real company or product name or you can create an unusual business name.
+This will make the name of your business name consistent throughout your website.
+2, Decide on your business name.
+You will be choosing the official name of your business for all activities that represent your business.
+This includes your business name, products, your website, and any …
+30% Building a website can be done in 10 simple steps: 1.
+Preliminaries and planning 2.
+Designing 3.
+Prototyping 4.
+Production 5.
+Operation… Building a website can be done in 10 simple steps: To: With A Web You Are Into 2 Zim 40 Wh I M T If Y As An Onl You Orwice There W Os You Hcimp …
+40% Building a website can be done in 10 simple steps: 2 days of website plan.
+If you want to set a web site.
+This is the 1st step.
+In the 2 days .
+It wants the company to develop a website to help the investors and investors.
+He will have another investor.
+He wants to help others.
+The company is working to grow, and create a name for some investors and investors.
+100 million US dollars It is used to build a hotel and a hotel.
+In order to obtain their cooperation, investors, investors, investors in .
+Building a website can be done in 10 simple steps: Front 1222ED825333 $3333$etennenn eenneenneSecondinjustanyanojunatajin …
+A.5 More Experiments on compressing different LLMs
+Table 11: Zero-shot performance comparison of OPT-6.7B, LLaMA 2-7B, Vicuna-7B, Mistral-7B compressed by SVD-LLM and baselines under 20% compression ratio on seven common sense reasoning datasets (measured by both individual and average accuracy ($\uparrow$)).
+The best performance is marked in bold.
+The relative performance gain compared to the best-performing baseline is marked in green color inside bracket.
+[…]
+OPT-6.7B SVD 0.14 0.27 0.51 0.25 0.23 0.54 0.21 0.31
+FWSVD 0.15 0.26 0.49 0.26 0.20 0.52 0.21 0.30
+ASVD 0.14 0.40 0.51 0.30 0.20 0.59 0.22 0.34
+SVD-LLM 0.24 0.60 0.60 0.45 0.28 0.73 0.24 0.45 ($\uparrow$32%)
+LLaMA 2-7B SVD 0.15 0.27 0.49 0.26 0.23 0.52 0.20 0.30
+FWSVD 0.12 0.25 0.49 0.25 0.22 0.52 0.21 0.30
+ASVD 0.25 0.31 0.60 0.41 0.32 0.72 0.23 0.41
+SVD-LLM 0.26 0.50 0.60 0.41 0.26 0.66 0.23 0.41 ($\uparrow$0%)
+Mistral-7B SVD 0.14 0.25 0.52 0.26 0.23 0.54 0.20 0.30
+FWSVD 0.15 0.28 0.52 0.26 0.21 0.53 0.21 0.31
+ASVD 0.21 0.51 0.58 0.42 0.25 0.50 0.26 0.39
+SVD-LLM 0.17 0.55 0.58 0.36 0.25 0.67 0.21 0.40 ($\uparrow$3%)
+Vicuna-7B SVD 0.15 0.26 0.50 0.26 0.23 0.52 0.20 0.30
+FWSVD 0.14 0.27 0.49 0.26 0.22 0.53 0.20 0.30
+ASVD 0.21 0.53 0.55 0.39 0.30 0.41 0.23 0.37
+SVD-LLM 0.23 0.51 0.58 0.40 0.28 0.67 0.22 0.41 ($\uparrow$11%)
+We also evaluate the performance of different LLMs, including OPT-6.7B, LLaMA 2-7B, Mistral-7B, and Vicuna-7B under 20% compression ratio on seven common sense reasoning datasets.
+The results are shown in Section A.5.
+SVD-LLM performs better than the best-performing baseline in most of the datasets across different LLMs and even achieves 32% higher average accuracy on OPT-6.7B.